### v2

# TabTreeFormer: Tabular Data Generation Using Hybrid Tree-Transformer

Abstract
Transformers have achieved remarkable success in tabular data generation. However, they lack domain-specific inductive biases which are critical to preserving the intrinsic characteristics of tabular data. Meanwhile, they suffer from poor scalability and efficiency due to quadratic computational complexity. In this paper, we propose TabTreeFormer, a hybrid transformer architecture that incorporates a tree-based model that retains tabular-specific inductive biases of non-smooth and potentially low-correlated patterns due to its discreteness and non-rotational invariance, and hence enhances the fidelity and utility of synthetic data. In addition, we devise a dual-quantization tokenizer to capture the multimodal continuous distribution and further facilitate the learning of numerical value distribution. Moreover, our proposed tokenizer reduces the vocabulary size and sequence length due to the limited dimension-wise semantic meaning and training set size of tabular data, rendering a significant model size shrink without sacrificing the capability of the transformer model. We evaluate TabTreeFormer on 10 datasets against multiple generative models on various metrics; our experimental results show that TabTreeFormer achieves superior fidelity, utility, privacy, and efficiency. Our best model yields a 40% utility improvement with 1/16 of the baseline model size.
Tabular Data, Generative Models
1 Introduction
Figure 1: Performance comparison of TabTreeFormer (Ours) and other representative tabular generative models in terms of fidelity, utility (i.e., MLE), privacy, and efficiency. TabTreeFormer achieves the best overall performance and outperforms other models on almost all sole metrics.
Tables are one of the most prevalent data modalities in real-world applications (healthcare (Dash et al., 2019), financial services (Assefa et al., 2021), etc.), yet are heavily under-exploited due to privacy concerns (EU, 2016). Fortunately, synthetic data offers an alternative to the utilization of tabular data by modeling the characteristics of real data and reducing the risk of data breach (PDPC, 2024), thus drawing considerable attention in recent years.
The state-of-the-art research shows that transformers such as autoregressive transformers (Borisov et al., 2023), masked transformers (Gulati & Roysdon, 2024), and diffusion models with transformers (Zhang et al., 2024) have achieved impressive fidelity, utility, and privacy in tabular data generation, enabling synthetic data to empower a variety of fields (Hernandez et al., 2022; Assefa et al., 2021). However, unlike the application of transformers in other research areas such as computer vision (Dosovitskiy et al., 2021) and natural language processing (Vaswani et al., 2017), existing transformer models for synthetic tabular data often overlook domain-specific priors (i.e., inductive biases). For example, CvT (Wu et al., 2021) introduces convolutional embedding and projection to transformers to boost vision performances, and Transformer-XL (Dai et al., 2019) introduces segment-level recurrence to transformers to capture longer-term dependencies in long text. While vision and language models have enjoyed the performance boost introduced by domain-specific priors, less exploration has been conducted to leverage them in tabular generative models. Moreover, transformers suffer from poor scalability due to quadratic computational complexity. This raises two interesting questions:
  1. 1.
    Are tabular-specific inductive biases beneficial to the performance of tabular generative models?
  2. 2.
    Can one exploit unique characteristics of tabular data to improve model scalability and efficiency?
To answer the questions, we propose TabTreeFormer, a generative hybrid transformer that incorporates a tree-based model and a deliberately devised tokenizer to process and introduce tabular-specific inductive biases. We also leverage the limited per-dimension semantic meaning and limited training set size of typical tabular data to enable drastically reduced model size by our tokenizer design.
Inspired by the superior performance of tree-based models on tabular tasks (Shwartz-Ziv & Armon, 2022; Gorishniy et al., 2024), we propose to employ them on capturing tabular-specific inductive biases of non-smoothness and potentially low-correlation (Grinsztajn et al., 2024). Non-smoothness is attributed to the existence of discrete features and non-smooth relations between discrete and/or continuous features. Neural networks are biased to overly smooth solutions while tree-based models are exceptionally good at capturing the non-smooth patterns since they learn piece-wise constant functions rather than low-frequency functions (i.e., the functions neural networks tend to bias toward) (Rahaman et al., 2019). Low-correlated features includes uninformative features that contribute trivially to downstream tasks such as classification and regression. Due to the non-rotationally invariant nature of tree-based models (Ng, 2004), trees are more robust against low-correlated features, while neural networks are biased towards the universal presence of correlation.
Nevertheless, trees do not capture multimodal distribution in continuous features, which is a tabular inductive bias particularly essential for generative tasks (Xu et al., 2019; Shi et al., 2024). We then propose dual-quantization tokenization of numeric values, where the first quantization is a K-Means (MacQueen, 1967) quantization that captures multimodal distribution, and the second quantization is a quantile quantization that is used to present the precise value.
Considering the structured nature of tabular data and limited semantic meaning in each dimension, we further optimize the tokenizer so that the vocabulary size and sequence length can be reduced significantly. Together with the limited of training data compared to vision or language tasks, this leads to room for reducing other hyperparameters such as hidden dimensions, and eventually the entire model size.
Our contributions are summarized as follows:
  • To the best of our knowledge, we are the first work that introduces tabular-specific inductive biases to improve the performance of tabular generative models. We propose a novel hybrid TFM coupled with a TBM to capture such inductive biases.
  • We devise a dual-quantization tokenizer that models the multimodal distribution of continuous values in real data. This practice facilitates the learning of the numerical distribution and further improves the quality of synthetic tabular data.
  • Employing the characteristics of tabular data, our tokenizer design also leads to a much smaller vocabulary size and sequence length, achieving a phenomenal model size reduction and efficiency enhancement.
  • We evaluate TabTreeFormer across 10 datasets against multiple representative tabular generative models and show that TabTreeFormer outperforms other models in fidelity, utility, privacy, and efficiency, as shown in Fig. 1. Notably, with just 1/16 of the baseline model size, our best model achieves a 40% utility improvement over the best-performing baseline model.
2 Related Works
2.1 Tabular Data Generation
Early works on tabular data generation predominantly utilized MLPs and CNNs as backbone architectures and GAN and VAE as generation methods (Park et al., 2018; Xu et al., 2019; Zhao et al., 2021, 2024). Recent works evolved to use transformers with auto-regression, masked modeling, and diffusion as generative paradigms (Borisov et al., 2023; Zhang et al., 2024; Gulati & Roysdon, 2024). For example, TabMT (Gulati & Roysdon, 2024) employs a masked transformer with ordered embedding and achieves outstanding utility in downstream tasks with improved scalability, and TabSyn (Zhang et al., 2024) encodes tabular data into a latent space and generates tabular data using diffusion model, demonstrating impressive fidelity and utility. Despite their enhanced performance, these works overlook the essential inductive biases such as non-smoothness and low-correlated features that are crucial to preserving the intrinsic characteristics of tabular data.
Xu et al. (2019) finds that multimodal distribution in continuous features is a critical inductive bias for table generation. CTGAN (Xu et al., 2019) uses variational Gaussian mixture model (Bishop, 2006) to decompose multimodal values while TabDiff (Shi et al., 2024) introduces a multimodal stochastic sampler. In this work, we introduce multimodal distribution, into an auto-regressive transformer, which was absent in prior works.
2.2 Tree-based Models for Tabular Data
In recent years, tree-based gradient boosting algorithms have emerged as the dominant approach for tabular data tasks, particularly with the introduction of models like XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018). These tree-based models are ensembles of small and intuitive decision trees that naturally understand feature relations and interactions, thus achieving high predictive performance, efficiency, and scalability, making them the preferred choice in both academia and industry. While some recent research has shown promise of transformers in tabular data (Huang et al., 2020; Arik & Pfister, 2021; Hollmann et al., 2023) in certain scenarios, their computational cost and suboptimal performance make them still less favorable compared to tree-based models (Shwartz-Ziv & Armon, 2022; Gorishniy et al., 2024), which is attributed to tabular-specific inductive biases, including non-smoothness and potentially low-correlation, captured by tree-based models (Grinsztajn et al., 2024). In this work, we leverage these features of tree-based models and retain tabular-specific inductive bias to facilitate the training of our transformer model.
2.3 Tokenizers for Tabular Data
Transformers are not originally designed for tabular data, entailing a need for a tokenizer to translate tabular data to transformer-friendly input tokens or embeddings. Arik & Pfister (2021); Zhang et al. (2024) translate tabular data into a continuous space by completely reformulating the tokenization and embedding layers. Borisov et al. (2023); Solatorio & Dupriez (2023); Gulati & Roysdon (2024) translate tabular data to a sequence of tokens (i.e., discrete space) by a tabular-specific tokenization method, and retain the embedding layer with moderate modifications, which is akin to natural language and more suitable for auto-regressive and masked transformers. In this paper, we opt for the second choice, to optimize the tokenizer for tabular data.
3 TabTreeFormer
Figure 2: Overview of TabTreeFormer. Tokens representing tree leaves are prepended to each row’s tokenized representation, which is also heavily optimized. Tree leaves and optimized data tokens introduce essential inductive biases to the tabular data generator.
Recall that our goal is to improve the performance and efficiency of the generative model by introducing domain-specific inductive biases and shrinking the model size. Formally, let $\mathbf{X}$ be the tabular training dataset for a generative model with $n$ rows, $m_{d}$ discrete and $m_{c}$ continuous features, we aim to generate synthetic tabular data $\mathbf{X}^{\prime}$ similar to $\mathbf{X}$. To achieve this, we propose TabTreeFormer that consists of three building blocks: a tree-based model, a tokenizer, and a transformer model, as shown in Fig. 2. The tree-based model captures some tabular-specific inductive biases, and the tokenizer is used to grasp the multimodal distribution while reducing vocabulary size and sequence length simultaneously. The transformer model learns the prior extracted from the tree-based model and tokenizer and produces high-quality synthetic data. Concrete accounts of training and generation algorithms are described in Algorithm 1 and 2.
Algorithm 1 Training of TabTreeFormer
1:  Input: Tabular dataset $\mathbf{X}$ ($n$ rows, $m_{d}$ discrete features, $m_{c}$ continuous features), $s$ total training steps
2:  Output: Fitted tree-based model $\mathcal{T}$ and the leaf index matrix $\mathbf{J}$, data tokenizer $\mathcal{D}$, and transformer-based language model $\mathcal{G}$
[…]
8:  $\mathcal{G}\leftarrow\textsc{InitializeLM}(\text{n\_pos}=2+T+m_{d}+2m_{c},\text{n\_vocab}=3+n_{l}+n_{c}+n_{b}+n_{q})$
9:  for $i=1,\dots,s$ do
10:     $\mathbf{B}\leftarrow\textsc{SampleBatch}(\mathbf{Z})$
11:     $\widetilde{\mathbf{B}}\leftarrow\textsc{Mask}(\mathbf{B})$
12:     $\mathcal{G}\leftarrow\textsc{UpdateGradient}(\text{input}=\widetilde{\mathbf{B}},\text{target}=\mathbf{B})$
13:  end for
Algorithm 2 Tabular data generation using TabTreeFormer
1:  Input: Leaf index matrix $\mathbf{J}$, data tokenizer $\mathcal{D}$, transformer-based language model $\mathcal{G}$, and number of rows to be generated $n^{\prime}$
2:  Output: Synthetic tabular dataset $\mathbf{X}^{\prime}$
3:  $\mathbf{J}^{\prime}\leftarrow\textsc{Sample}(\mathbf{J},n^{\prime})$
4:  $\mathbf{Z}^{\prime}\leftarrow\textsc{Mask}(\textsc{ToTokenIDs}(\mathbf{J}^{\prime}))$
5:  $\widetilde{\mathbf{X}}^{\prime}\leftarrow\mathcal{G}.\textsc{Generate}(\text{partial\_input}=\mathbf{Z}^{\prime})$
6:  $\mathbf{X}^{\prime}\leftarrow\mathcal{D}.\textsc{Decode}(\widetilde{\mathbf{X}}^{\prime})$
3.1 Tree-based Models
To introduce tabular-specific inductive biases, we incorporate tree-based models into the transformer (“Tree-based Model” in Fig. 2). This is achieved by augmenting each token sequence representing one table row with additional information derived from the leaf indices of a tree-based model. Formally, let the fitted tree-based model be $\mathcal{T}$ with $T$ trees. Let $l_{i}\in\mathbb{N}$ be the number of leaves in $i$-th tree, where $i\in\{1,2,\dots,T\}$. On $\mathcal{T}$, we have the leaf index matrix $\mathbf{J}=[j_{ik}]\in\mathbb{N}^{n\times T}$ of $\mathbf{X}$ where $j_{ik}\in\{1,2,\dots,l_{k}\}$ represents the leaf index of $i$-th row $\mathbf{X}_{i}$ in the $k$-th tree. Leaf indices provide flattened positional information of each row in a tree, and hence inherently introduce tabular inductive biases captured by trees – non-smoothness and potentially low-correlations. With leaf indices prepended to actual data values, although the subsequent step is still a neural network, tabular inductive biases from tree-based models are inherited by the entire synthetic data generation model through the construction of the intermediate input sequence of tokens, thereby improving the quality of the generated synthetic data. Leaf indices are also provided as prompts during generation, allowing indirect yet informative hints to generate more realistic data.
3.2 Tokenizer
To model the multimodal distribution of continuous values and reduce tokenized value sizes to eventually reduce the model size, we devise a tailored data tokenizer for tabular data (“Tokenizer” in Fig. 2). For each continuous feature, we quantize the values so that it is of a more natural data format for transformers to learn. However, unlike TabMT (Gulati & Roysdon, 2024) that uses a single quantizer, we use a dual-quantizer (i.e., double quantizers) to capture multimodal continuous distribution and describe the precise values respectively. Transformers do not expect near-Gaussian-distributed input tokens, so we capture the multimodal distribution simply by a K-Means quantizer (MacQueen, 1967) with a small number of clusters (e.g. $K=10$). Following that, the original value is then quantized into a large number of quantiles (e.g. $Q=1000$), based on which original values are recovered, and thus the values are highly precise. By the dual-quantization tokenization method, each numeric value are encoded into 2 discrete values: a bin ID in the K-Means quantizer, and a quantile ID in the quantile quantizer, making 2 tokens sufficient to represent a numeric value with useful and sufficient information. Categorical features are simply tokenized by label encoding, requiring 1 token. Formally, a reversible data tokenizer $\mathcal{D}$ can be fitted on $\mathbf{X}$. Let $c_{i}\in\mathbb{N},b_{i}\in\{0,1,\dots,K\},q_{i}\in\{0,1,\dots,Q\}$ denote the number of categories, K-Means bins, and quantiles per column respectively, where $c_{i}=0$ in continuous features and $b_{i}=q_{i}=0$ in categorical features. Then, the domain for a categorical value is $\{1,2,\dots,c_{i}\}\subseteq\mathbb{N}_{+}$, and the domain for a continuous value is $\{1,2,\dots,b_{i}\}\times\{1,2,\dots,q_{i}\}\subseteq\mathbb{N}_{+}^{2}$.
3.3 Auto-regressive Transformer
Table 1: Illustration of tokens in TabTreeFormer. “Size” indicates the number of tokens belonging to this type. Recall Figure 2 for the colors and usage of them. “N” in “Format” column can be any index in the range. For example, leaf 3 ($3<n_{l}$) is thus represented as [leaf3], corresponding to the light green tokens in Figure 2 with value 3.
Type (Alias) Size Format Source Description
Leaf $n_{l}=\max_{i\in\{1,2,\dots,T\}}l_{i}$ [Uncaptioned image] Tree-based model $\mathcal{T}$ Leaf index in a tree
Categorical (cat) $n_{c}=\max_{i\in\{1,2,\dots,n\}}c_{i}$ [Uncaptioned image] Data tokenizer $\mathcal{D}$ (categorical columns) Category ID
Cluster (bin) $n_{b}=\max_{i\in\{1,2,\dots,n\}}b_{i}$ [Uncaptioned image] Data tokenizer $\mathcal{D}$ (numeric columns) Bin ID in K-Means quantizer
Quantile (quant) $n_{q}=\max_{i\in\{1,2,\dots,n\}}q_{i}$ [Uncaptioned image] Data tokenizer $\mathcal{D}$ (numeric columns) Quantile ID in quantile quantizer
Special 3 [Uncaptioned image] Classical special tokens for LMs BOS, EOS, and mask tokens
TabTreeFormer is simply trained as an auto-regressive transformer. The input token sequence is constructed by concatenating the outcome from the tree model and tokenizer. Totally 5 major types of tokens are required and the vocabulary size is $V=n_{l}+n_{c}+n_{b}+n_{q}+3$. The number of tokens in a constructed sequence of tokens corresponding to one row is $L=2+T+m_{d}+2m_{c}$, including BOS and EOS tokens. Both of them are much lower than the need for natural language models. A shared set of leaf tokens across all trees and similarly shared sets of categorical, bin, and quantile tokens across all features are used. Distinctions between different trees and features are effectively encoded through positional information, leveraging the positional embeddings inherent in most transformers.
A key observation on the construction of token sequence is that the valid range of tokens at each position is fixed for a given table. Consequently, tokens can be sampled exclusively from this precomputed set of valid options during generation. This ensures that every generated token sequence corresponds to a valid row, eliminating the need for rejecting invalid samples – a common practice in existing methods (Borisov et al., 2023; Solatorio & Dupriez, 2023; Zhao et al., 2023) – and significantly accelerating the sampling process.
Additionally, transformers are prone to memorization, particularly when trained on tabular data, where datasets are much smaller compared to typical natural language corpora (Borisov et al., 2023). To avoid over-fitting in terms of memorization, we heavily mask the input and maintain the target output unmasked. The rest of the training and generation settings are similar to the training and generation of classical causal language models, except for a modified loss function to account for the ordinal relations between quantile tokens, with details introduced in Section 3.4.
3.4 Loss Function
We use cross-entropy loss (CEL) as the optimization objective of the transformer. However, the quantized quantile tokens retain an inherent ordinal property, where closer IDs correspond to closer values. This relationship is not captured by standard CEL, which treats all classes as equally distinct. To address this, we replace the vanilla CEL with a specialized ordinal CEL for quantile tokens.
Formally, let the predicted logits for a token be $\mathbf{z}\in\mathbb{R}^{V}$, and let $t\in\{1,\dots,V\}$ be the target label. Recall the classical definition of CEL as given in Equation 1.
$\mathcal{L}_{\text{ce}}(\mathbf{z},t)=-\log\frac{e^{z_{t}}}{\sum_{i=1}^{V}e^{z_{i}}}$ (1)
We define ordinal cross-entropy loss (OCEL) as a variant of CEL where each term is weighted. These weights are applied to unnormalized probabilities in both the numerator and denominator, depending on the current- and target-class pairs. It contrasts with the classical weighted CEL, which weights the loss per class based solely on the target class and applies to the numerator only. Formally, we write it as Equation 2, where $w_{ti}$ denotes the weight for $z_{i}$, where $t$ is the target class.
$\mathcal{L}_{\text{oce}}(\mathbf{z},t)=-\log\frac{w_{tt}e^{z_{t}}}{\sum_{i=1}^{V}w_{ti}e^{z_{i}}}$ (2)
Lemma 1.
Given $\mathbf{w}>\mathbf{0}$ not dependent on $\mathbf{z}$, OCEL is optimized when $z_{t}\to\infty$ and $z_{i}\to-\infty,\forall i\neq t$.
Theorem 1.
If $w_{ti}\geq 0$ is a function of the distance of the current class $i$ to the target class $t$, namely, $w_{ti}=f(|t-i|)$ where $f:\mathbb{N}\rightarrow\mathbb{R}_{+},|t-i|\mapsto w_{i}$, and $f$ is monotonically increasing, then the OCEL penalizes adjacent classes less than farther classes.
Both Lemma 1 and Theorem 1 are intuitive, and we provide rigorous proofs in Appendix A.1-A.2.
Theorem 1 requires the definition of $w_{ti}=f(|t-i|)$ that satisfies the monotonically increasing constraint to be used for OCEL. Any satisfying definition works technically. In this paper, we define the weight $w_{ti}$ for $z_{i}$ to be obtained by Equation 3.
$w_{ti}=1+m-e^{-\frac{(t-i)^{2}}{(V\sigma)^{2}}}\vspace{-0.5em}$ (3)
where $\sigma=0.005$ is a scaling factor of the distance, and $m=0.5$ is the minimum weight (used on $i=t$). Detailed explanation and analysis of it can be found in Appendix A.3.
Note that not all tokens are part of the ordinal relation, so the OCEL is applied only to valid quantile tokens at the corresponding positions. To enable the model to learn to generate valid quantile tokens, we introduce an additional modified CEL to distinguish valid tokens from invalid ones, which complements OCEL. Formally, let valid quantile tokens have IDs in the range $[Q_{s},Q_{e})$ The valid token group loss (for quantile tokens at this position) is defined by Equation 4 when $t\in[Q_{s},Q_{e})$.
$\mathcal{L}_{\text{vg}}(\mathbf{z},t)=-\log\frac{\sum_{i\in[Q_{s},Q_{e})}e^{z_{i}}}{\sum_{i=1}^{V}e^{z_{i}}}\vspace{-0.5em}$ (4)
In sum, Equation 3.4 expresses the loss function on one token.
$\mathcal{L}(\mathbf{z},t)$
$=$ $\begin{cases}\mathcal{L}_{\text{ce}}(\mathbf{z},t)&\text{ if }t\notin[Q_{s},Q_{e})\\ \mathcal{L}_{\text{oce}}(\mathbf{z}_{Q_{s}:Q_{e}},t-Q_{s})+\mathcal{L}_{\text{vg}}(\mathbf{z},t)&\text{ if }t\in[Q_{s},Q_{e})\\ \end{cases}\vspace{-0.8em}$ (5)
Although the ordinal property also holds for K-Means bin IDs, we do not calculate the loss on them this way because the number of different bins is small, and we expect the model to predict the bins accurately.
4 Experiments
4.1 Experimental Setup
We conduct all experiments on NVIDIA RTX 4090. We use LightGBM (Ke et al., 2017) whose hyperparameters are tuned by Optuna (Akiba et al., 2019) as the tree-based model. We exploit Distill-GPT2 (Sanh et al., 2019) as the transformer backbone, following the practice in prior works (Borisov et al., 2023; Solatorio & Dupriez, 2023). TabTreeFormer (TTF) has three configurations: S (small), M (medium), and L (large), with the number of trainable parameters of approximately 1M, 5M, and 40M, respectively. Implementation and setting details can be found in Appendix B and Appendix C.1.
Datasets. Experiments are conducted on 10 datasets with diverse sizes and characteristics from OpenML (Vanschoren et al., 2013): adult, bank, boston, breast, cahouse, credit, diabetes, iris, qsar, and wdbc. Details of these datasets are summarized in Appendix C.2.
Baseline models. We compare the performance of TabTreeFormer against several influential and high-performing methods in tabular data generation including non-neural networks, GANs, VAEs, diffusion models, and auto-regressive transformers: ARF (Watson et al., 2023), CTAB-GAN+ (CTAB+) (Zhao et al., 2024), CTGAN (Xu et al., 2019), TVAE (Xu et al., 2019), TabDDPM (Kotelnikov et al., 2023), TabSyn (Zhang et al., 2024), and GReaT (Borisov et al., 2023), REaLTabFormer (RTF) (Solatorio & Dupriez, 2023). Implementation of baselines is described in Appendix C.3.
Metrics. We comprehensively evaluate TabTreeFormer using fidelity, utility, privacy, and efficiency, which are standard metrics in tabular data generation (Xu et al., 2019; Zhao et al., 2024; Borisov et al., 2023; Zhang et al., 2024; Shi et al., 2024). Detailed implementation of the metrics can be found in Appendix C.4-C.7.
  • Fidelity shows the cosmetic discrepancy between the real and synthetic data. It is evaluated via “Shape” and “Trend” metrics  (Dat, 2023; Shi et al., 2024). “Shape” measures the similarity of marginal distribution density for each column, and “Trend” measures the fidelity in correlation between column pairs. Higher “Shape” and “Trend” values indicate better data fidelity.
  • Utility exhibits the quality of synthetic data by testing its performance on downstream tasks. We evaluate based on the established Train-on-Synthetic, Test-on-Real (TSTR) machine learning efficacy (MLE) evaluation framework (Xu et al., 2019), namely, synthetic data generated based on real training data is tested on a hold-out real test set. Three models are used: linear (i.e. linear and logistic regression; LN), random forest (RF) (Breiman, 2001), and XGBoost (XGB) (Chen & Guestrin, 2016). Better generative models tend to yield better scores on downstream tasks.
  • Privacy is a crucial criterion when the synthetic data serves the use case of privacy-preserving data sharing. We use the distance to the closest record (DCR) (Zhao et al., 2021) to evaluate the privacy of synthetic data. We compare the DCR from synthetically generated data and from hold-out real (test) data to the real training data and privacy preservation is demonstrated by DCRs calculated on the latter being no smaller than the former, tested by Mann-Whitney U Test (Mann & Whitney, 1947). Privacy is at risk of disclosure when $p$-values are smaller than 0.05.
  • Efficiency showcases the model sizes and computation time for training and generation of different models. Under comparable performance, smaller models and faster computation are favored.
4.2 Fidelity
Figure 3: Performance comparison between different models in terms of fidelity. The higher the values are, the better the fidelity is. Top-4 performing baselines by fidelity mean scores are visualized.
Figure 4: Visualization of the marginal densities of the generated data of baseline tabular auto-regressive transformers and TabTreeFormer in comparison to the real data on some representative multimodal-distributed continuous columns. All these models use Distill-GPT2 (or its structure and hyperparameters) as backbone.
Figure 5: Performance comparison of TabTreeFormer versus top-4 performing baselines in fidelity on pair-wise correlation in 3 experimented datasets. Real (absolute) correlation values are presented on the left, and the absolute error in correlation values in synthetic data from different models, capped at 0.4 for better visibility, are shown at right (the darker, the better). The left-top features are categorical fatures, and right-bottom features are numeric features.
Figure 6: Real vs. synthetic abs. corr.
Figure 7: Feature count vs. “Trend” score improvement of TabTreeFormer from some baselines.
Marginal density distribution of columns (Shape). TabTreeFormer achieves comparable performance in “Shape” metric, as shown in Fig. 3. Meanwhile, it can be seen from Fig. 4 that TabTreeFormer captures better multimodal distribution compared to other auto-regressive transformers, which validates the effectiveness of our dual-quantization tokenizer design. More fidelity results can be found in Table 9 and Fig. 11 in Appendix D.1.
Pair-wise correlation between columns (Trend). TabTreeFormer outperforms all baselines in “Trend” metric with a significant margin, as presented in Fig. 3, implying the superior capability of TabTreeFormer in learning inter-feature relations. It can be observed from Fig. 7-7 that even strong baselines struggle with correlations involving categorical features and correlation values are not close to 0. Moreover, TabTreeFormer tends to show a more significant improvement from auto-regressive transformer baselines on datasets with a larger number of features as referred in Fig. 7. This also demonstrates the advantage of introducing the inductive bias on low-correlations so that the models learn relations among a larger number of features better by filtering out less correlated features.
4.3 Utility (MLE)
Table 2: MLE performance between different models. The top two rows calculate the global average scores and relative error (w.r.t real data). The middle three rows show the average performance in each ML model, and the bottom two rows show the average performances in classification (clf.) and regression (reg.). Classification tasks are evaluated by weighted AUC ROC scores, and regression tasks are evaluated by $R^{2}$ scores. Best scores and second best scores are highlighted in bold with and without underscore, respectively.
ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
Avg. $0.843_{\pm 0.155}$ $0.680_{\pm 0.271}$ $0.810_{\pm 0.189}$ $0.552_{\pm 0.865}$ $0.800_{\pm 0.206}$ $0.866_{\pm 0.130}$ $0.865_{\pm 0.123}$ $0.813_{\pm 0.189}$ $0.863_{\pm 0.127}$ $\boldsymbol{0.878}_{\pm 0.113}$ $\underline{\boldsymbol{0.880}}_{\pm 0.111}$
RE $0.057_{\pm 0.077}$ $0.244_{\pm 0.274}$ $0.099_{\pm 0.137}$ $0.483_{\pm 1.214}$ $0.110_{\pm 0.170}$ $0.027_{\pm 0.040}$ $0.027_{\pm 0.029}$ $0.095_{\pm 0.142}$ $0.029_{\pm 0.040}$ $\boldsymbol{0.017}_{\pm 0.023}$ $\underline{\boldsymbol{0.016}}_{\pm 0.023}$
LN $0.835_{\pm 0.176}$ $0.686_{\pm 0.258}$ $0.810_{\pm 0.193}$ $0.570_{\pm 0.882}$ $0.800_{\pm 0.197}$ $0.851_{\pm 0.153}$ $0.849_{\pm 0.156}$ $0.814_{\pm 0.163}$ $0.853_{\pm 0.154}$ $\boldsymbol{0.861}_{\pm 0.145}$ $\underline{\boldsymbol{0.862}}_{\pm 0.143}$
RF $0.842_{\pm 0.156}$ $0.681_{\pm 0.273}$ $0.806_{\pm 0.200}$ $0.566_{\pm 0.910}$ $0.798_{\pm 0.220}$ $0.873_{\pm 0.122}$ $0.872_{\pm 0.109}$ $0.810_{\pm 0.210}$ $0.866_{\pm 0.121}$ $\boldsymbol{0.888}_{\pm 0.095}$ $\underline{\boldsymbol{0.891}}_{\pm 0.090}$
XGB $0.851_{\pm 0.147}$ $0.672_{\pm 0.307}$ $0.814_{\pm 0.194}$ $0.519_{\pm 0.897}$ $0.802_{\pm 0.222}$ $0.874_{\pm 0.125}$ $0.874_{\pm 0.110}$ $0.814_{\pm 0.209}$ $0.870_{\pm 0.116}$ $\boldsymbol{0.884}_{\pm 0.102}$ $\underline{\boldsymbol{0.886}}_{\pm 0.101}$
clf. $0.907_{\pm 0.074}$ $0.750_{\pm 0.227}$ $0.895_{\pm 0.065}$ $0.850_{\pm 0.114}$ $0.868_{\pm 0.130}$ $0.917_{\pm 0.064}$ $0.913_{\pm 0.071}$ $0.890_{\pm 0.081}$ $0.918_{\pm 0.062}$ $\underline{\boldsymbol{0.922}}_{\pm 0.063}$ $\boldsymbol{0.922}_{\pm 0.065}$
reg. $0.586_{\pm 0.122}$ $0.400_{\pm 0.266}$ $0.472_{\pm 0.127}$ $-0.641_{\pm 1.467}$ $0.528_{\pm 0.239}$ $0.660_{\pm 0.122}$ $0.674_{\pm 0.095}$ $0.506_{\pm 0.187}$ $0.643_{\pm 0.061}$ $\boldsymbol{0.698}_{\pm 0.083}$ $\underline{\boldsymbol{0.710}}_{\pm 0.092}$
The MLE results are presented in Table 2, whose raw scores are in Table 10 in Appendix D.2, show that TabTreeFormer outperforms baseline models steadily regardless of the complexity of the ML model used, task, and dataset. TabTreeFormer-S achieves similar performance as the best-performing baselines (REaLTabForme and TabSyn), and the M and L versions of TabTreeFormer outperforms them by a significant margin of around 40%, assessed by relative errors of ML performance compared to real data. Furthermore, out of the 30 MLE scores, TabTreeFormer-M or -L outperforms all other baselines in 20 cases (2/3), and they both outperform all other baselines in 16 cases ($\geq$1/2).
As an important foundation of the superior capability of tree-based models in capturing potentially low-correlations inductive bias, rotational invariance of models is an aspect worthwhile analyzing. As visualized in Fig. 9, rotationally invariant generators (e.g. CTAB-GAN+, CTGAN, GReaT, TVAE) are generally worse than non-rotationally invariant counterparts (e.g. REaLTabFormer and our TabTreeFormer by positional embedding in a fixed order and TabSyn by VAE-encoding of data). Also, TabTreeFormer’s design outperforms rotationally invariant baseline generators more on better-performing non-rotationally invariant downstream models (RF and XGB) than on rotationally invariant downstream models (LN). This again emphasizes the importance of applying non-rotationally invariant models to introduce tabular inductive bias (for low correlations).
Figure 8: MLE relative error to real data with TabTreeFormer-L. On the left of the red line are rotationally invariant models.
Figure 9: Number of features vs. average generation time per (valid) row.
4.4 Privacy (DCR)
Table 3: Summarized performance comparison between different models in terms of privacy. The first row shows the number of violations, and the second row shows the minimum $p$ value for privacy leakage risk, the closer the value to 0, the higher the risk of privacy leakage. Risky values are highlighted in red.
ARF GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
# vio. 0 5 0 0 1 1 0 0
min. $p$ 0.071 0.000 0.097 0.141 0.003 0.045 0.061 0.111
As summarized in Table 3 (raw scores in Table 11 in Appendix D.3), TabTreeFormer-M and L are safe in all the 10 datasets. TabTreeFormer-S has a privacy leakage in only one dataset, with the $p$-value very close to 0.05. In comparison, GReaT has severe privacy concerns in 5 out of 10 datasets. Therefore, while auto-regressive transformers have privacy leakage concerns, TabTreeFormer and REaLTabFormer are generally privacy-preserving, validating the effectiveness of using masking in preserving privacy for tabular auto-regressive transformers.
4.5 Efficiency
Table 4: Relative performance change of TabTreeFormer-M compared to TabTreeFormer-L.
Fidelity Utility
Shape $-0.208\%$ Avg. MLE scores $-0.225\%$
Trend $-0.436\%$ Avg. MLE relative error $-9.624\%$
Observing results reported above, the performance of TabTreeFormer-M is better than all baselines in most metrics, while it has only about 5M parameters and its most competitive counterparts have more than 20M (TabSyn) or even more than 80M parameters (GReaT and REaLTabFormer). Also, as shown in Table 4, there is no significant performance difference between TabTreeFormer-M and TabTreeFormer-L, although their sizes differ by a factor of 8. This demonstrates the capability of TabTreeFormer in maintaining high efficiency without significant compromise in effectiveness.
Table 5: Training time (s) comparison of TabTreeFormer compared to other auto-regressive transformer baselines on large datasets. Best values are highlighted in bold with underscore, and second best values are highlighted in bold without underscore.
Dataset adult bank cahouse
GReaT $10397.040_{\pm 9.203}$ $10295.769_{\pm 21.729}$ $4373.950_{\pm 56.089}$
RTF $1875.904_{\pm 34.663}$ $835.167_{\pm 1.619}$ $\boldsymbol{596.523}_{\pm 51.206}$
TTF-M $\underline{\boldsymbol{235.554}}_{\pm 71.542}$ $\underline{\boldsymbol{225.707}}_{\pm 39.181}$ $\underline{\boldsymbol{340.992}}_{\pm 76.571}$
TTF-L $\boldsymbol{606.790}_{\pm 277.828}$ $\boldsymbol{744.726}_{\pm 221.487}$ $946.903_{\pm 102.584}$
Compared to baseline auto-regressive transformers, TabTreeFormer demonstrates significant advantage in terms of efficiency, especially in training time on large datasets (see Table 5) and in generation time on datasets with a large number of features (see Fig. 9), which is consistent with our expectation due to the smaller model size and sampling without rejection. Raw training and generation time records can be found in Table 13 in Appendix D.5.
4.6 Ablation Study
Table 6: Ablation studies. The values show the relative score change when certain component in TabTreeFormer-S is taken out.
LN RF XGB
w/o TBM $-4.121\%$ $-4.798\%$ $-7.657\%$
w/o DQT $-6.141\%$ $-0.297\%$ $-0.273\%$
w/o OCEL $+0.452\%$ $-0.967\%$ $-1.446\%$
On each core components of TabTreeFormer, including tree-based model (TBM), dual-quantization tokenization (DQT), and OCEL, we apply ablation study using TabTreeFormer-S on 2 randomly selected datasets: credit and diabetes, evaluated using MLE. The results, presented in Table 6 with raw values presented in Table 14 in Appendix D.6, verify the usefulness of all the core components.
5 Conclusion
In this paper, we propose TabTreeFormer, a hybrid tree-transformer for generating high-quality tabular data. TabTreeFormer inherits unique inductive biases from a tree-based model and also captures multimodal distribution to optimize the data representations of numeric values via a dual-quantization tokenizer. We comprehensively evaluate the performance of TabTreeFormer and show that it demonstrates superior fidelity, utility, privacy and outstanding efficiency with significantly smaller model size and computation time.
Appendix A Supplementary Loss Function Explanation and Analysis
A.1 Proof of Lemma 1
Proof.
Firstly, note that $\frac{\partial w_{ti}}{\partial z_{j}}=0,\forall i,j\in\{1,\dots,V\}$, by independence between the two. Also mind that $w_{ti}>0,\forall i\in\{1,\dots,V\}$.
$\frac{\partial\mathcal{L}_{\text{oce}(\mathbf{z},t)}}{\partial z_{t}}$
$=$ $-\frac{\sum_{i=1}^{V}w_{ti}e^{z_{i}}}{w_{tt}e^{z_{t}}}\cdot\frac{w_{tt}e^{z_{t}}\sum_{i=1}^{V}w_{ti}e^{z_{i}}-w_{tt}e^{z_{t}}\cdot w_{tt}e^{z_{t}}}{\left(\sum_{i=1}^{V}w_{ti}e^{z_{i}}\right)^{2}}$
$=$ $-\frac{\sum_{i=1}^{V}w_{ti}e^{z_{i}}-w_{tt}{e^{z_{t}}}}{\sum_{i=1}^{V}w_{ti}e^{z_{i}}}<0$
$\frac{\partial\mathcal{L}_{\text{oce}(\mathbf{z},t)}}{\partial z_{i}}$
$=$ $-\frac{\sum_{i=1}^{V}w_{ti}e^{z_{i}}}{w_{tt}e^{z_{t}}}\cdot\frac{-w_{ti}e^{z_{i}}\cdot w_{tt}e^{z_{t}}}{\left(\sum_{i=1}^{V}w_{ti}e^{z_{i}}\right)^{2}}$
$=$ $\frac{w_{ti}e^{z_{i}}}{\sum_{i=1}^{V}w_{ti}e^{z_{i}}}>0$
Thus, $\mathcal{L}_{\text{oce}}(\mathbf{z},t)$ is optimized (minimized) when $z_{t}\to\infty,z_{i}\to-\infty(t\neq i)$. ∎
A.2 Proof of Theorem 1
Proof.
Suppose we have $\mathbf{z}$ and $\mathbf{z}^{*}$ as logits on vocabulary size $V$ and target label $t$, which satisfy both the following constraints:
  • For some $i\neq j\in\{1,\dots,V\}\setminus$ and $|t-i|\neq|t-j|$, $a=z_{i}=z_{j}^{*}\neq z_{j}=z_{i}^{*}=b$ for some $a,b\in\mathbb{R}$.
  • $z_{k}=z_{k}^{*},\forall k\in\{0,1,\dots,V-1\}\setminus\{i,j,t\}$.
Without loss of generality, suppose $d_{1}=|t-i|<|t-j|=d_{2},a<b$. Note that $w_{k}$ is monotonically increasing with $|t-k|$.
Thus, $e^{a}<e^{b},w_{ti}<w_{tj}$. This gives $\left(w_{ti}-w_{tj}\right)\left(e^{a}-e^{b}\right)<0$, so $P=w_{ti}e^{z_{i}}+w_{tj}e^{z_{j}}=w_{ti}e^{a}+w_{tj}e^{b}<w_{ti}e^{z_{i}^{*}}+w_{tj}e^{z_{j}^{*}}=w_{ti}e^{b}+w_{tj}e^{a}=Q$ for some $P,Q\in\mathbb{R}$.
Substitute into $\mathcal{L}_{\text{oce}}(\mathbf{z},t)$ and $\mathcal{L}_{\text{oce}}(\mathbf{z}^{*},t)$,
$\mathcal{L}_{\text{oce}}(\mathbf{z},t)$ $=-\log\frac{w_{tt}e^{z_{t}}}{P+\sum_{k\neq i,j}w_{k}e^{z_{k}}}$
$\mathcal{L}_{\text{oce}}(\mathbf{z}^{*},t)$ $=-\log\frac{w_{tt}e^{z_{t}}}{Q+\sum_{k\neq i,j}w_{k}e^{z_{k}}}$
Hence, $\mathcal{L}_{\text{oce}}(\mathbf{z},t)<\mathcal{L}_{\text{oce}}(\mathbf{z}^{*},t)$. The logits with higher logit value at position closer to target token is penalized less. ∎
A.3 Detailed Explanations and Analysis of the Weight Function (Equation 3)
Recall the function of weight given the target token and current token in Equation 3. This section will provide detailed explanations and analysis of it.
The exponential term comes from the intuition that the difference between closer quantiles is much more significant than the difference between farther quantiles. Some scaling and shifting is applied on the exponential term.
Figure 10 illustrates the relationship between $w_{ti}$ and $i$ under varying values of $\sigma$ and $m$. The parameter $\sigma$ primarily governs the width around the target token where the weights are significantly altered; larger $\sigma$ values result in a broader width. In contrast, $m$ dictates the intensity of the weight modification, with smaller $m$ values leading to more pronounced changes in weight.
Figure 10: The relation of weights $w_{ti}$ (normalized so that $\|\mathbf{w}\|=1$ for visualization purpose to token index $i$, with vocabulary size $V=1000$, and target token $t=350$.
Appendix B TabTreeFormer Model Setup Details
B.1 Tree-based Model
B.1.1 Hyperparameter Tuning
For the tree-based classifier or regressor, we use Optuna (Akiba et al., 2019) to tune the hyperparameters of LightGBM (Ke et al., 2017) using the training data. If the dataset has a designated target column, the model is trained to predict that column. Otherwise, a random column can be selected as the target. Then, fit the model with the selected hyperparameters for TabTreeFormer.
The hyperparameter space explored is
  • Learning Rate: Logarithmic scale float in $[0.01,0.3]$;
  • Number of Estimators: Integers in $[50,250]$ step 50;
  • Maximum Depth: Integers in $[3,10]$;
  • Number of Leaves: Integers in $[20,100]$ step 5;
  • Minimum Size per Leaf: Integers in $[10,50]$ step 5;
  • Feature Fraction: Float in $[0.6,1.0]$;
  • Bagging Fraction: Float in $[0.6,1.0]$;
  • L1 Regularization: Float in $[0.0,10.0]$;
  • L2 Regularization: Float in $[0.0,10.0]$.
Optimization was performed over 50 trials using 3-fold cross-validation, performance evaluated based on weighted F1 for classification and (negative) mean squared error for regression.
B.1.2 Leaf Index Matrix
In the subsequent steps of TabTreeFormer, leaf index matrix $\mathbf{J}$ of the real data $\mathbf{X}$ is computed. The indexing order of trees and their leaves used in the subsequent steps are directly taken from the fitted tree-based model. The leaf index matrix is essentially the result of .apply function for most tree-based models in sklearn.
B.2 Tokenization
For K-Means and quantile quantization, we use sklearn.preprocessing.KBinsDiscretizer with strategy “kmeans” and “quantile” respectively.
In cases where the number of distinct values expected to be kept is way larger than typical natural language models’ vocabulary size, two or more consecutive quantile quantizers may be used, where subsequent quantizers are used to quantize the values in the same quantile as suggested by previous quantizers. Thus, with $q$ quantizers, a total of $Q^{q}$ values can be represented, which is sufficient to cover most practical cases. In this paper, we present $Q=1000$ as sufficient precision for simplicity.
For categorical values, we use sklearn.preprocessing.OrdinalEncoder.
B.3 Transformer
B.3.1 Model Configuration
For the language model, we use Distill-GPT2 (Sanh et al., 2019), based on the GPT architecture (Radford et al., 2019), with several modifications. The vocabulary size and maximum sequence length are adjusted based on the dataset, and the values are likely much lower than the requirement for natural language models. For example, if $n_{l}=200,n_{c}=20,n_{b}=K=10,n_{q}=Q=1000$, then $V=1233$, and if $T=200,m_{d}=10,m_{c}=10$, then $L=222$, while natural language models typically have $V>30000,L>1000$.
In this paper, we present 3 versions of TabTreeFormer of different sizes. Table 7 shows their details. The model sizes are obtained for diabetes (Smith et al., 1988) dataset. Actual values may vary based on the fitted tree-based model and dataset.
Table 7: Hyperparameter configuration of TabTreeFormer of different sizes.
TabTreeFormer S M L
Approx. #Params 1 5 40
hidden state dimensions 128 256 768
inner feed-forwared layer dimensions 256 1024 3072
number of attention heads per layer 8 8 12
B.3.2 Token Sequence Construction
The tokens in Table 1 are conceptual tokens. When converting the leaf, category, K-Means bin, and quantile IDs into token IDs in actual implementation, we do not explicitly make the textual tokens, but instead, directly add corresponding token type’s offset.
B.3.3 Masking
Instead of using a fixed mask ratio, we apply varying mask ratios to enhance the diversity of training data. Tree-related tokens (yellow section in Figure 2) are masked with a ratio sampled uniformly from $[0.3,0.7]$, while actual value tokens (blue section in Figure 2) are masked with a ratio sampled uniformly from $[0.1,0.4]$.
To ensure that the quantile token is always masked if its corresponding K-Means bin token is masked within the same numeric column, we first generate a random mask. If a violation is detected (i.e., the K-Means bin token is unmasked while the quantile token is masked), we swap the mask values of these two positions. This approach ensures that masks are generated both efficiently and accurately.
B.3.4 Training Configuration
The reduced model size allows for a significantly larger batch size compared to typical natural language models. We use a batch size of 128 per device, reduce by half in case of CUDA OOM [^1] Among all $10\times 3=30$ experiments, TabTreeFormer-L hits CUDA OOM in one experiment, and no other TabTreeFormer experiments had memory issues with this batch size., and train for up to 3,000 steps with a learning rate of $5\times 10^{-4}$ in FP16 precision, utilizing the Hugging Face Trainer.
B.3.5 Model Inference
Tree-leaf tokens from real data after masking are used as prompts for generation. We apply a temperature of 0.7 with sampling for generation.
Recall that a set of valid tokens can be pre-determined at each position. Formally, let $p_{i}$ be the corresponding feature index for the $i$-th token from data tokenizer, and $R_{i}$ indicates the type of this token. For example, in Figure 2, $p_{1}=1,p_{2}=1,p_{3}=2,R_{1}=\text{bin},R_{2}=\text{quant},R_{3}=\text{cat}$. Then, let $\mathcal{V}_{i}$ be the set of tokens allowed at position $i\in\{1,2,\dots,L\}$ in the entire sequence, we have Equation 6.
$\mathcal{V}_{i}=\begin{cases}\{\text{[BOS]}\}&\text{ if }i=1\\ \{\text{[leaf1],...,[leaf(}l_{i-1}\text{)]}&\text{ if }2\leq i\leq T+1\\ \{\text{[bin1],...,[bin(}b_{p_{i-T-1}}\text{)]}&\text{ if }T+2\leq i\leq L-1\text{ and }R_{i-T-1}=\text{bin}\\ \{\text{[quant1],...,[quant(}q_{p_{i-T-1}}\text{)]}&\text{ if }T+2\leq i\leq L-1\text{ and }R_{i-T-1}=\text{quant}\\ \{\text{[cat1],...,[cat(}b_{p_{i-T-1}}\text{)]}&\text{ if }T+2\leq i\leq L-1\text{ and }R_{i-T-1}=\text{cat}\\ \{\text{[EOS]}\}&\text{ if }i=L\end{cases}$
Valid tokens for a specific position is enforced by setting the logits of invalid tokens at the position to negative infinity before sampling.
Appendix C Experiment Setup Details
C.1 Experiment Environment
Experiments were conducted on a system running Ubuntu 22.04.1 LTS with kernel version 6.8.0. The machine featured a 13th Gen Intel Core i9-13900K CPU with 24 cores and 32 threads, capable of a maximum clock speed of 5.8 GHz. GPU computations were performed using a single NVIDIA GeForce RTX 4090 with 24 GB VRAM. The environment was set up with NVIDIA driver version 565.57.01 and CUDA 12.7.
C.2 Details on Datasets
Table 8 shows the details of the datasets we used. Datasets are downloaded by sklearn.datasets.fetch_openml, with the dataset name as input.
Table 8: Datasets used in experiments. #R, #F, #N, and #C represents the number of rows, features (including target column), numeric features, and categorical features respectively. The task can be classification, denoted “clf” followed by the number of classes in the bracket, and regression, denoted “reg”. Aliases in bracket in “Name” column show the names used in this paper.
Name #R #F #N #C Task
adult (Kohavi, 1996) 48842 15 2 13 clf(2)
bank-marketing (bank) (Moro et al., 2011) 45211 17 7 10 clf(2)
boston (Harrison & Rubinfeld, 1978) 506 14 12 2 reg
breast-w (breast) (Wolberg & Mangasarian, 1990) 699 10 9 1 clf(2)
california_housing (cahouse) (Kelley Pace & Barry, 1997) 20640 10 9 1 reg
credit-g (credit) (Hofmann, 1994) 1000 21 7 14 clf(2)
diabetes (Smith et al., 1988) 768 9 8 1 clf(2)
iris (Fisher, 1936) 150 5 4 1 clf(3)
qsar-biodeg (qsar) (Mansouri et al., 2013) 1055 42 41 1 clf(2)
wdbc (Street et al., 1993) 569 31 30 1 clf(2)
C.3 Baseline Implementation
We use the Synthcity (Qian et al., 2024) implementation for ARF (Watson et al., 2023), CTGAN (Xu et al., 2019), GReaT (Borisov et al., 2023), TabDDPM (Kotelnikov et al., 2023), and TVAE (Xu et al., 2019). We also include CTAB-GAN+ (Zhao et al., 2024), REaLTabFormer (Solatorio & Dupriez, 2023), and TabSyn (Zhang et al., 2024) using their official code on GitHub or public SDK.
C.4 Fidelity Metrics Implementation
Fidelity metrics are calculated using SDMetrics (Dat, 2023)’s public SDK.
C.5 Machine Learning Efficacy (MLE) Metrics Implementation
Instead of running all models with fixed, typically default, parameters across all datasets, we perform hyperparameter tuning before reporting their performance. The tuning process follows the same methodology used to optimize the core tree-based model in TabTreeFormer, but with 30 trials per round of optimization.
For linear models, no hyperparameter tuning is performed, as they follow the classical definition without additional tunable parameters.
The hyperparameter space explored for random forest is
  • Number of Estimators: Integers in $[100,300]$ step 50;
  • Maximum Depth: Integers in $[5,20]$ step 5;
  • Minimum Samples Split: Integers in $[2,10]$ step 2;
  • Minimum Size per Leaf: Integers $[1,5]$;
  • Maximum Features: Value in “sqrt”, “log2”, and NULL;
  • Bootstramp: Enabled or disabled.
The hyperparameter space explored for XGBoost is:
  • Learning Rate: Logarithmic scale float in $[0.01,0.3]$;
  • Number of Estimators: Integers in $[100,300]$ step 50;
  • Maximum Depth: Integers in $[3,10]$;
  • Minimum Child Weight: Logarithmic scale float in $[1.0,10.0]$;
  • Minimum Split Loss Gamma: Float in $[0.0,0.5]$;
  • Subsample Ratio: Float in $[0.5,1.0]$;
  • Subsample Ratio of Columns by Tree: Float in $[0.5,1.0]$;
  • L1 Regularization: Float in $[0.0,10.0]$.
For all models, numeric values are standardized using standard scaling. Categorical values are one-hot encoded for linear models and label-encoded for random forest and XGBoost. To simplify the experiments, rows with missing values are excluded, as handling missing data is not the focus of this paper.
The reported performance for machine learning utility is measured using the weighted AUC-ROC for classification tasks and the $R^{2}$ score for regression tasks, so for all these scores, a larger value indicates a better performance. Each generator is trained and used to generate synthetic data of the same size as the training dataset three times for each dataset, and the summarized results of these runs are reported.
C.6 Distance to Closest Record (DCR) Metrics Implementation
The way the DCR values are reported differ from paper to paper. Also, existing papers often report a value and claim some specific value to be an ideal value, overlooking the fact that better privacy usually sacrifices quality. Thus, in this paper, we instead apply a statistical test on DCR values to validate privacy-preserving capability of the models, instead of evaluating it, which is likely also more useful in practical privacy-preserving data sharing cases.
We compare the DCR to the real training dataset $\mathbf{X}$ of a hold-out real dataset $\widehat{\mathbf{X}}$ with a synthetic dataset of the same size $\mathbf{X}^{\prime}$, and privacy-perserving means that DCRs calculated on the latter is not smaller than DCRs calculated on the former. We test this using Mann-Whitney U Test (Mann & Whitney, 1947), with the null hypothesis $H_{0}$ being that the distance between $\widehat{\mathbf{X}}$ and $\mathbf{X}$ is greater than or equal to $\mathbf{X}^{\prime}$ and $\mathbf{X}$, and if the $p$-value is less than 0.05, $\mathbf{X}^{\prime}$ is closer to $\mathbf{X}$ than $\widehat{\mathbf{X}}$, implying a risk of privacy leakage.
The data is preprocessed by standard scaling for numeric values and one-hot encoding for categorical values for distance calculation. Cosine distance (calculated using sklearn) is used to evaluate the distance between records. As a reference, we calculate the cosine distances between the real test set and the real train set, obtaining the minimum distance per record for the test set. Similarly, we compute the cosine distances between synthetic data (with the same number of rows as the real test set) and the real train set, and also obtain the minimum distances. These distances are compared with the reference values to assess the similarity and privacy-preserving characteristics of the synthetic data.
C.7 Real Data Detection (RDD) Metrics Implementation
This metric is not discussed in the main paper as it does not directly pertain to the primary use cases of synthetic data. However, since it is a standard evaluation framework, we include it in the appendix for completeness.
The RDD metric evaluates how easily real data can be distinguished from synthetic data. A higher difficulty in detection indicates better synthetic data quality. The evaluation is framed as a binary classification task, with the experimental settings kept the same as those used for machine learning efficacy (MLE). However, unlike MLE, where higher scores are better, for RDD, lower AUC-ROC values indicate better performance, as they reflect greater similarity between synthetic and real data.
The training data for the RDD task consists of the real training set combined with a synthetic dataset of the same size as the real training set. The test data comprises the real test set combined with another synthetic dataset of the same size as the real test set.
Appendix D Supplementary Experimental Results
D.1 Supplementary Fidelity Results
Table 9: Performance comparison between different models in terms of fidelity. The closer the values are to 1, the better the fidelity is. Best scores and second best scores are highlighted in bold with and without underscore, respectively. Equal values in the first 3 digits may be compared by the 4-th or 5-th digits.
Dataset Metric ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
adult Shape $0.942_{\pm 0.001}$ $0.964_{\pm 0.005}$ $0.881_{\pm 0.008}$ $\underline{\boldsymbol{0.983}}_{\pm 0.001}$ $0.929_{\pm 0.000}$ $0.962_{\pm 0.001}$ $\boldsymbol{0.977}_{\pm 0.011}$ $0.877_{\pm 0.004}$ $0.946_{\pm 0.002}$ $0.954_{\pm 0.001}$ $0.956_{\pm 0.001}$
Trend $0.876_{\pm 0.004}$ $0.915_{\pm 0.004}$ $0.760_{\pm 0.005}$ $\boldsymbol{0.953}_{\pm 0.006}$ $0.882_{\pm 0.000}$ $0.933_{\pm 0.003}$ $\underline{\boldsymbol{0.955}}_{\pm 0.018}$ $0.726_{\pm 0.002}$ $0.869_{\pm 0.017}$ $0.883_{\pm 0.020}$ $0.899_{\pm 0.001}$
bank Shape $0.947_{\pm 0.000}$ $0.965_{\pm 0.002}$ $0.891_{\pm 0.001}$ $\underline{\boldsymbol{0.983}}_{\pm 0.001}$ $0.915_{\pm 0.000}$ $0.967_{\pm 0.001}$ $\boldsymbol{0.983}_{\pm 0.005}$ $0.894_{\pm 0.003}$ $0.906_{\pm 0.006}$ $0.929_{\pm 0.001}$ $0.937_{\pm 0.001}$
Trend $0.885_{\pm 0.019}$ $0.884_{\pm 0.003}$ $0.870_{\pm 0.009}$ $0.932_{\pm 0.030}$ $0.903_{\pm 0.000}$ $0.955_{\pm 0.001}$ $\underline{\boldsymbol{0.967}}_{\pm 0.008}$ $0.870_{\pm 0.011}$ $0.925_{\pm 0.011}$ $0.954_{\pm 0.001}$ $\boldsymbol{0.962}_{\pm 0.002}$
boston Shape $0.907_{\pm 0.005}$ $0.850_{\pm 0.003}$ $0.795_{\pm 0.023}$ $0.489_{\pm 0.025}$ $\boldsymbol{0.916}_{\pm 0.000}$ $\underline{\boldsymbol{0.942}}_{\pm 0.004}$ $0.893_{\pm 0.003}$ $0.801_{\pm 0.009}$ $0.850_{\pm 0.001}$ $0.855_{\pm 0.000}$ $0.856_{\pm 0.001}$
Trend $0.921_{\pm 0.005}$ $0.828_{\pm 0.011}$ $0.896_{\pm 0.004}$ $0.606_{\pm 0.006}$ $0.908_{\pm 0.000}$ $0.947_{\pm 0.002}$ $0.934_{\pm 0.003}$ $0.843_{\pm 0.005}$ $0.944_{\pm 0.003}$ $\boldsymbol{0.968}_{\pm 0.001}$ $\underline{\boldsymbol{0.969}}_{\pm 0.000}$
breast Shape $\boldsymbol{0.854}_{\pm 0.005}$ $0.775_{\pm 0.024}$ $0.832_{\pm 0.019}$ $0.838_{\pm 0.001}$ $0.728_{\pm 0.000}$ $0.753_{\pm 0.014}$ $0.813_{\pm 0.005}$ $\underline{\boldsymbol{0.900}}_{\pm 0.013}$ $0.820_{\pm 0.011}$ $0.851_{\pm 0.002}$ $0.852_{\pm 0.008}$
Trend $0.767_{\pm 0.006}$ $0.616_{\pm 0.063}$ $0.709_{\pm 0.015}$ $0.767_{\pm 0.000}$ $0.636_{\pm 0.000}$ $0.668_{\pm 0.010}$ $0.709_{\pm 0.006}$ $0.781_{\pm 0.010}$ $0.749_{\pm 0.010}$ $\underline{\boldsymbol{0.799}}_{\pm 0.003}$ $\boldsymbol{0.796}_{\pm 0.016}$
cahouse Shape $\boldsymbol{0.979}_{\pm 0.001}$ $0.962_{\pm 0.003}$ $0.908_{\pm 0.005}$ $0.974_{\pm 0.006}$ $0.879_{\pm 0.000}$ $0.974_{\pm 0.003}$ $\underline{\boldsymbol{0.991}}_{\pm 0.001}$ $0.944_{\pm 0.001}$ $0.949_{\pm 0.006}$ $0.975_{\pm 0.001}$ $0.979_{\pm 0.000}$
Trend $0.971_{\pm 0.002}$ $0.927_{\pm 0.002}$ $0.914_{\pm 0.004}$ $0.946_{\pm 0.002}$ $0.913_{\pm 0.000}$ $0.969_{\pm 0.003}$ $\underline{\boldsymbol{0.989}}_{\pm 0.002}$ $0.930_{\pm 0.003}$ $0.947_{\pm 0.013}$ $0.969_{\pm 0.002}$ $\boldsymbol{0.974}_{\pm 0.000}$
credit Shape $0.971_{\pm 0.003}$ $0.937_{\pm 0.013}$ $0.944_{\pm 0.003}$ $0.926_{\pm 0.007}$ $0.932_{\pm 0.000}$ $0.944_{\pm 0.003}$ $0.932_{\pm 0.007}$ $0.931_{\pm 0.004}$ $0.963_{\pm 0.001}$ $\boldsymbol{0.979}_{\pm 0.001}$ $\underline{\boldsymbol{0.983}}_{\pm 0.000}$
Trend $0.925_{\pm 0.004}$ $0.862_{\pm 0.030}$ $0.894_{\pm 0.005}$ $0.864_{\pm 0.007}$ $0.861_{\pm 0.000}$ $0.896_{\pm 0.006}$ $0.862_{\pm 0.008}$ $0.848_{\pm 0.004}$ $0.903_{\pm 0.002}$ $\boldsymbol{0.933}_{\pm 0.003}$ $\underline{\boldsymbol{0.946}}_{\pm 0.000}$
diabetes Shape $0.898_{\pm 0.004}$ $0.923_{\pm 0.006}$ $0.793_{\pm 0.042}$ $0.769_{\pm 0.017}$ $0.881_{\pm 0.000}$ $\boldsymbol{0.943}_{\pm 0.004}$ $\underline{\boldsymbol{0.955}}_{\pm 0.007}$ $0.824_{\pm 0.012}$ $0.934_{\pm 0.003}$ $0.939_{\pm 0.002}$ $0.943_{\pm 0.002}$
Trend $0.949_{\pm 0.007}$ $0.898_{\pm 0.004}$ $0.913_{\pm 0.001}$ $0.875_{\pm 0.003}$ $0.921_{\pm 0.000}$ $0.952_{\pm 0.005}$ $0.964_{\pm 0.002}$ $0.858_{\pm 0.008}$ $0.965_{\pm 0.004}$ $\boldsymbol{0.967}_{\pm 0.001}$ $\underline{\boldsymbol{0.970}}_{\pm 0.004}$
iris Shape $0.893_{\pm 0.011}$ $0.822_{\pm 0.022}$ $0.762_{\pm 0.040}$ $0.553_{\pm 0.024}$ $0.870_{\pm 0.000}$ $\underline{\boldsymbol{0.906}}_{\pm 0.017}$ $\boldsymbol{0.893}_{\pm 0.003}$ $0.809_{\pm 0.004}$ $0.868_{\pm 0.007}$ $0.868_{\pm 0.000}$ $0.862_{\pm 0.006}$
Trend $0.897_{\pm 0.007}$ $0.590_{\pm 0.019}$ $0.793_{\pm 0.047}$ $0.506_{\pm 0.036}$ $0.899_{\pm 0.000}$ $0.896_{\pm 0.016}$ $0.917_{\pm 0.007}$ $0.783_{\pm 0.017}$ $0.931_{\pm 0.004}$ $\underline{\boldsymbol{0.942}}_{\pm 0.005}$ $\boldsymbol{0.934}_{\pm 0.002}$
qsar Shape $0.931_{\pm 0.004}$ $0.917_{\pm 0.003}$ $0.891_{\pm 0.001}$ $0.602_{\pm 0.005}$ $0.887_{\pm 0.000}$ $\underline{\boldsymbol{0.959}}_{\pm 0.003}$ $0.930_{\pm 0.005}$ $0.873_{\pm 0.007}$ $0.940_{\pm 0.001}$ $0.947_{\pm 0.000}$ $\boldsymbol{0.948}_{\pm 0.000}$
Trend $0.911_{\pm 0.009}$ $0.863_{\pm 0.007}$ $0.887_{\pm 0.010}$ $0.605_{\pm 0.001}$ $0.855_{\pm 0.000}$ $0.890_{\pm 0.007}$ $0.914_{\pm 0.009}$ $0.849_{\pm 0.012}$ $0.931_{\pm 0.008}$ $\boldsymbol{0.947}_{\pm 0.001}$ $\underline{\boldsymbol{0.949}}_{\pm 0.001}$
wdbc Shape $0.923_{\pm 0.003}$ $0.861_{\pm 0.005}$ $0.824_{\pm 0.031}$ $0.395_{\pm 0.017}$ $0.873_{\pm 0.000}$ $0.945_{\pm 0.006}$ $0.946_{\pm 0.004}$ $0.749_{\pm 0.026}$ $0.951_{\pm 0.001}$ $\underline{\boldsymbol{0.975}}_{\pm 0.001}$ $\boldsymbol{0.972}_{\pm 0.003}$
Trend $0.939_{\pm 0.001}$ $0.866_{\pm 0.007}$ $0.934_{\pm 0.003}$ $0.772_{\pm 0.010}$ $0.896_{\pm 0.000}$ $0.954_{\pm 0.001}$ $0.976_{\pm 0.002}$ $0.907_{\pm 0.017}$ $0.963_{\pm 0.003}$ $\underline{\boldsymbol{0.986}}_{\pm 0.001}$ $\boldsymbol{0.985}_{\pm 0.001}$
Avg. Shape $0.925_{\pm 0.038}$ $0.898_{\pm 0.067}$ $0.852_{\pm 0.059}$ $0.751_{\pm 0.224}$ $0.881_{\pm 0.059}$ $\boldsymbol{0.929}_{\pm 0.065}$ $\underline{\boldsymbol{0.931}}_{\pm 0.054}$ $0.860_{\pm 0.062}$ $0.913_{\pm 0.050}$ $0.927_{\pm 0.051}$ $0.929_{\pm 0.052}$
Trend $0.904_{\pm 0.056}$ $0.825_{\pm 0.120}$ $0.857_{\pm 0.076}$ $0.783_{\pm 0.161}$ $0.867_{\pm 0.084}$ $0.906_{\pm 0.089}$ $0.919_{\pm 0.083}$ $0.839_{\pm 0.061}$ $0.913_{\pm 0.064}$ $\boldsymbol{0.935}_{\pm 0.055}$ $\underline{\boldsymbol{0.938}}_{\pm 0.056}$
The raw experiment results for fidelity is shown in Table 9.
Figure 11: Visualization of the marginal densities of the generated data of baseline auto-regressive transformers and TabTreeFormer in comparison to the real data on some more representative multimodal-distributed continuous columns.
Figure 11 shows the performance of TabTreeFormer in comparison to other auto-regressive transformers in capturing multimodal distributions. TabTreeFormer shows a better capability in capturing such distribution in general.
D.2 Supplementary Utility Results
The raw experiment results for MLE is shown in Table 10. The results are also visualized in Figure 12.
Table 10: Performance comparison between different models in terms of MLE. Classification tasks are evaluated by weighted AUC ROC scores, and regression tasks are evaluated by $R^{2}$ scores. The last row shows the average relative error versus real. Best scores and second best scores are highlighted in bold with and without underscore, respectively. Equal values in the first 3 digits may be compared by the 4-th or 5-th digits.
Dataset ML real ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
adult LN 0.914 $0.907_{\pm 0.000}$ $0.904_{\pm 0.001}$ $0.886_{\pm 0.009}$ $0.908_{\pm 0.000}$ $\boldsymbol{0.911}_{\pm 0.000}$ $\underline{\boldsymbol{0.911}}_{\pm 0.000}$ $0.910_{\pm 0.002}$ $0.893_{\pm 0.001}$ $0.905_{\pm 0.001}$ $0.908_{\pm 0.002}$ $0.910_{\pm 0.000}$
RF 0.910 $0.903_{\pm 0.001}$ $0.902_{\pm 0.002}$ $0.884_{\pm 0.003}$ $0.904_{\pm 0.001}$ $0.905_{\pm 0.000}$ $\underline{\boldsymbol{0.908}}_{\pm 0.000}$ $0.905_{\pm 0.003}$ $0.892_{\pm 0.001}$ $0.900_{\pm 0.000}$ $0.905_{\pm 0.000}$ $\boldsymbol{0.906}_{\pm 0.001}$
XGB 0.914 $0.907_{\pm 0.001}$ $0.905_{\pm 0.001}$ $0.886_{\pm 0.002}$ $0.908_{\pm 0.000}$ $0.911_{\pm 0.000}$ $\underline{\boldsymbol{0.912}}_{\pm 0.001}$ $0.910_{\pm 0.003}$ $0.895_{\pm 0.001}$ $0.904_{\pm 0.000}$ $0.909_{\pm 0.001}$ $\boldsymbol{0.911}_{\pm 0.001}$
bank LN 0.907 $0.895_{\pm 0.002}$ $0.901_{\pm 0.003}$ $0.887_{\pm 0.003}$ $0.900_{\pm 0.002}$ $\underline{\boldsymbol{0.907}}_{\pm 0.000}$ $0.904_{\pm 0.000}$ $0.900_{\pm 0.005}$ $0.890_{\pm 0.005}$ $0.894_{\pm 0.005}$ $\boldsymbol{0.905}_{\pm 0.001}$ $0.902_{\pm 0.001}$
RF 0.930 $0.900_{\pm 0.001}$ $0.902_{\pm 0.002}$ $0.882_{\pm 0.002}$ $0.908_{\pm 0.001}$ $0.907_{\pm 0.001}$ $0.913_{\pm 0.001}$ $0.907_{\pm 0.006}$ $0.884_{\pm 0.002}$ $0.893_{\pm 0.011}$ $\boldsymbol{0.916}_{\pm 0.002}$ $\underline{\boldsymbol{0.921}}_{\pm 0.001}$
XGB 0.936 $0.903_{\pm 0.005}$ $0.906_{\pm 0.000}$ $0.884_{\pm 0.002}$ $0.912_{\pm 0.001}$ $0.910_{\pm 0.001}$ $0.918_{\pm 0.002}$ $0.913_{\pm 0.007}$ $0.885_{\pm 0.001}$ $0.890_{\pm 0.012}$ $\boldsymbol{0.919}_{\pm 0.000}$ $\underline{\boldsymbol{0.927}}_{\pm 0.003}$
boston LN 0.590 $0.451_{\pm 0.021}$ $0.264_{\pm 0.034}$ $0.374_{\pm 0.075}$ $0.000_{\pm 1.391}$ $0.397_{\pm 0.000}$ $0.529_{\pm 0.028}$ $0.533_{\pm 0.036}$ $0.480_{\pm 0.030}$ $0.541_{\pm 0.024}$ $\boldsymbol{0.573}_{\pm 0.001}$ $\underline{\boldsymbol{0.577}}_{\pm 0.002}$
RF 0.662 $0.484_{\pm 0.057}$ $0.129_{\pm 0.015}$ $0.336_{\pm 0.089}$ $0.000_{\pm 1.288}$ $0.281_{\pm 0.001}$ $0.584_{\pm 0.006}$ $0.658_{\pm 0.025}$ $0.287_{\pm 0.032}$ $0.634_{\pm 0.028}$ $\boldsymbol{0.727}_{\pm 0.009}$ $\underline{\boldsymbol{0.738}}_{\pm 0.004}$
XGB 0.701 $0.509_{\pm 0.067}$ $0.097_{\pm 0.078}$ $0.364_{\pm 0.109}$ $0.000_{\pm 1.128}$ $0.277_{\pm 0.014}$ $0.578_{\pm 0.056}$ $0.647_{\pm 0.017}$ $0.284_{\pm 0.082}$ $0.653_{\pm 0.022}$ $\boldsymbol{0.677}_{\pm 0.023}$ $\underline{\boldsymbol{0.680}}_{\pm 0.003}$
breast LN 0.985 $0.984_{\pm 0.005}$ $0.913_{\pm 0.091}$ $0.978_{\pm 0.006}$ $0.985_{\pm 0.002}$ $0.985_{\pm 0.000}$ $\underline{\boldsymbol{0.987}}_{\pm 0.002}$ $\boldsymbol{0.987}_{\pm 0.001}$ $0.930_{\pm 0.009}$ $0.985_{\pm 0.003}$ $0.985_{\pm 0.001}$ $0.984_{\pm 0.001}$
RF 0.985 $0.982_{\pm 0.003}$ $0.968_{\pm 0.012}$ $\boldsymbol{0.985}_{\pm 0.003}$ $0.981_{\pm 0.002}$ $0.979_{\pm 0.004}$ $0.980_{\pm 0.000}$ $0.981_{\pm 0.002}$ $0.976_{\pm 0.007}$ $0.985_{\pm 0.002}$ $\underline{\boldsymbol{0.986}}_{\pm 0.001}$ $0.985_{\pm 0.001}$
XGB 0.984 $0.979_{\pm 0.001}$ $0.958_{\pm 0.025}$ $0.984_{\pm 0.002}$ $0.981_{\pm 0.001}$ $0.982_{\pm 0.002}$ $0.982_{\pm 0.002}$ $\underline{\boldsymbol{0.987}}_{\pm 0.001}$ $0.981_{\pm 0.005}$ $0.984_{\pm 0.002}$ $0.981_{\pm 0.000}$ $\boldsymbol{0.985}_{\pm 0.001}$
cahouse LN 0.653 $0.620_{\pm 0.003}$ $0.604_{\pm 0.012}$ $0.561_{\pm 0.019}$ $0.577_{\pm 0.013}$ $0.646_{\pm 0.000}$ $\underline{\boldsymbol{0.649}}_{\pm 0.001}$ $0.640_{\pm 0.002}$ $0.605_{\pm 0.002}$ $0.625_{\pm 0.004}$ $0.645_{\pm 0.005}$ $\boldsymbol{0.648}_{\pm 0.003}$
RF 0.822 $0.712_{\pm 0.010}$ $0.641_{\pm 0.015}$ $0.594_{\pm 0.016}$ $0.743_{\pm 0.002}$ $0.776_{\pm 0.003}$ $\boldsymbol{0.799}_{\pm 0.002}$ $0.771_{\pm 0.003}$ $0.679_{\pm 0.004}$ $0.700_{\pm 0.009}$ $0.774_{\pm 0.012}$ $\underline{\boldsymbol{0.801}}_{\pm 0.002}$
XGB 0.836 $0.737_{\pm 0.015}$ $0.664_{\pm 0.008}$ $0.605_{\pm 0.015}$ $0.769_{\pm 0.003}$ $0.792_{\pm 0.003}$ $\underline{\boldsymbol{0.819}}_{\pm 0.001}$ $0.792_{\pm 0.003}$ $0.700_{\pm 0.003}$ $0.708_{\pm 0.009}$ $0.792_{\pm 0.015}$ $\boldsymbol{0.814}_{\pm 0.003}$
credit LN 0.836 $0.773_{\pm 0.064}$ $0.538_{\pm 0.093}$ $0.814_{\pm 0.012}$ $0.781_{\pm 0.018}$ $0.665_{\pm 0.000}$ $0.801_{\pm 0.006}$ $0.780_{\pm 0.017}$ $0.702_{\pm 0.042}$ $0.823_{\pm 0.022}$ $\underline{\boldsymbol{0.823}}_{\pm 0.010}$ $\boldsymbol{0.822}_{\pm 0.006}$
RF 0.837 $0.769_{\pm 0.030}$ $0.505_{\pm 0.097}$ $0.782_{\pm 0.033}$ $0.755_{\pm 0.013}$ $0.722_{\pm 0.014}$ $\underline{\boldsymbol{0.820}}_{\pm 0.010}$ $0.791_{\pm 0.007}$ $0.746_{\pm 0.057}$ $0.811_{\pm 0.015}$ $\boldsymbol{0.796}_{\pm 0.008}$ $0.795_{\pm 0.004}$
XGB 0.844 $0.788_{\pm 0.019}$ $0.509_{\pm 0.088}$ $0.788_{\pm 0.023}$ $0.748_{\pm 0.026}$ $0.739_{\pm 0.006}$ $0.797_{\pm 0.039}$ $0.795_{\pm 0.014}$ $0.759_{\pm 0.048}$ $0.816_{\pm 0.015}$ $\underline{\boldsymbol{0.811}}_{\pm 0.016}$ $\boldsymbol{0.801}_{\pm 0.020}$
diabetes LN 0.884 $0.854_{\pm 0.014}$ $0.817_{\pm 0.028}$ $0.864_{\pm 0.017}$ $0.753_{\pm 0.032}$ $0.858_{\pm 0.000}$ $0.878_{\pm 0.003}$ $0.883_{\pm 0.006}$ $0.863_{\pm 0.008}$ $0.889_{\pm 0.002}$ $\boldsymbol{0.884}_{\pm 0.001}$ $\underline{\boldsymbol{0.885}}_{\pm 0.003}$
RF 0.869 $0.818_{\pm 0.023}$ $0.802_{\pm 0.014}$ $0.818_{\pm 0.008}$ $0.838_{\pm 0.024}$ $0.824_{\pm 0.012}$ $0.841_{\pm 0.010}$ $0.844_{\pm 0.012}$ $0.817_{\pm 0.020}$ $0.865_{\pm 0.009}$ $\underline{\boldsymbol{0.864}}_{\pm 0.006}$ $\boldsymbol{0.854}_{\pm 0.009}$
XGB 0.867 $0.838_{\pm 0.012}$ $0.811_{\pm 0.005}$ $0.807_{\pm 0.005}$ $0.832_{\pm 0.017}$ $0.831_{\pm 0.003}$ $0.849_{\pm 0.010}$ $0.847_{\pm 0.009}$ $0.820_{\pm 0.020}$ $0.863_{\pm 0.006}$ $\boldsymbol{0.849}_{\pm 0.008}$ $\underline{\boldsymbol{0.850}}_{\pm 0.012}$
iris LN 1.000 $0.997_{\pm 0.005}$ $0.272_{\pm 0.068}$ $0.886_{\pm 0.025}$ $0.983_{\pm 0.015}$ $0.985_{\pm 0.000}$ $0.983_{\pm 0.015}$ $0.999_{\pm 0.002}$ $0.937_{\pm 0.030}$ $0.997_{\pm 0.003}$ $\underline{\boldsymbol{1.000}}_{\pm 0.000}$ $\underline{\boldsymbol{1.000}}_{\pm 0.000}$
RF 1.000 $0.998_{\pm 0.002}$ $0.396_{\pm 0.099}$ $0.936_{\pm 0.018}$ $0.934_{\pm 0.058}$ $\underline{\boldsymbol{1.000}}_{\pm 0.000}$ $0.993_{\pm 0.005}$ $\underline{\boldsymbol{1.000}}_{\pm 0.000}$ $0.995_{\pm 0.004}$ $1.000_{\pm 0.000}$ $\underline{\boldsymbol{1.000}}_{\pm 0.000}$ $\underline{\boldsymbol{1.000}}_{\pm 0.000}$
XGB 1.000 $\underline{\boldsymbol{1.000}}_{\pm 0.000}$ $0.211_{\pm 0.042}$ $0.971_{\pm 0.009}$ $0.817_{\pm 0.203}$ $\underline{\boldsymbol{1.000}}_{\pm 0.000}$ $0.998_{\pm 0.003}$ $\underline{\boldsymbol{1.000}}_{\pm 0.000}$ $0.986_{\pm 0.004}$ $0.996_{\pm 0.004}$ $0.998_{\pm 0.003}$ $0.998_{\pm 0.002}$
qsar LN 0.906 $0.888_{\pm 0.011}$ $0.716_{\pm 0.085}$ $0.869_{\pm 0.019}$ $0.783_{\pm 0.057}$ $0.673_{\pm 0.000}$ $0.880_{\pm 0.003}$ $0.867_{\pm 0.012}$ $0.862_{\pm 0.015}$ $0.886_{\pm 0.005}$ $\underline{\boldsymbol{0.898}}_{\pm 0.010}$ $\boldsymbol{0.896}_{\pm 0.002}$
RF 0.936 $0.877_{\pm 0.021}$ $0.648_{\pm 0.036}$ $0.874_{\pm 0.020}$ $0.714_{\pm 0.080}$ $0.616_{\pm 0.009}$ $0.907_{\pm 0.007}$ $0.882_{\pm 0.006}$ $0.850_{\pm 0.026}$ $0.898_{\pm 0.010}$ $\boldsymbol{0.926}_{\pm 0.002}$ $\underline{\boldsymbol{0.927}}_{\pm 0.003}$
XGB 0.921 $0.872_{\pm 0.011}$ $0.731_{\pm 0.020}$ $0.874_{\pm 0.010}$ $0.713_{\pm 0.056}$ $0.602_{\pm 0.020}$ $0.897_{\pm 0.010}$ $0.860_{\pm 0.016}$ $0.856_{\pm 0.022}$ $0.895_{\pm 0.007}$ $\underline{\boldsymbol{0.913}}_{\pm 0.003}$ $\boldsymbol{0.907}_{\pm 0.002}$
wdbc LN 0.993 $0.979_{\pm 0.006}$ $0.929_{\pm 0.053}$ $0.983_{\pm 0.007}$ $0.946_{\pm 0.028}$ $0.979_{\pm 0.000}$ $0.988_{\pm 0.003}$ $0.992_{\pm 0.001}$ $0.979_{\pm 0.013}$ $0.990_{\pm 0.002}$ $\boldsymbol{0.992}_{\pm 0.001}$ $\underline{\boldsymbol{0.992}}_{\pm 0.001}$
RF 0.976 $0.972_{\pm 0.002}$ $0.919_{\pm 0.024}$ $0.974_{\pm 0.004}$ $0.894_{\pm 0.051}$ $0.976_{\pm 0.003}$ $0.980_{\pm 0.012}$ $\underline{\boldsymbol{0.984}}_{\pm 0.004}$ $0.976_{\pm 0.001}$ $0.975_{\pm 0.000}$ $\boldsymbol{0.981}_{\pm 0.002}$ $0.980_{\pm 0.001}$
XGB 0.990 $0.979_{\pm 0.001}$ $0.930_{\pm 0.021}$ $0.974_{\pm 0.008}$ $0.512_{\pm 0.082}$ $0.973_{\pm 0.003}$ $0.988_{\pm 0.002}$ $0.986_{\pm 0.002}$ $0.977_{\pm 0.006}$ $0.986_{\pm 0.002}$ $\underline{\boldsymbol{0.991}}_{\pm 0.001}$ $\boldsymbol{0.990}_{\pm 0.003}$
Avg. RE - $0.057_{\pm 0.077}$ $0.244_{\pm 0.274}$ $0.099_{\pm 0.137}$ $0.483_{\pm 1.214}$ $0.110_{\pm 0.170}$ $0.027_{\pm 0.040}$ $0.027_{\pm 0.029}$ $0.095_{\pm 0.142}$ $0.029_{\pm 0.040}$ $\boldsymbol{0.017}_{\pm 0.023}$ $\underline{\boldsymbol{0.016}}_{\pm 0.023}$
Figure 12: Performance comparison between different models in terms of MLE. Top-5 performing baselines by average MLE relative errors are visualized. LN, RF, and XGB are the ML model names.
D.3 Supplementary Privacy Results
The raw experiment results for DCR is shown in Table 11.
Table 11: Performance comparison between different models in terms of DCR. $p$-values less than $0.05$ is highlighted in red, indicating a high risk of privacy leakage.
Dataset ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
adult $0.486_{\pm 0.077}$ $0.691_{\pm 0.424}$ $0.687_{\pm 0.272}$ $0.440_{\pm 0.125}$ $0.000_{\pm 0.000}$ $0.139_{\pm 0.131}$ $0.654_{\pm 0.149}$ $0.003_{\pm 0.004}$ $0.999_{\pm 0.001}$ $0.992_{\pm 0.011}$ $0.993_{\pm 0.006}$
bank $0.999_{\pm 0.000}$ $0.849_{\pm 0.093}$ $0.996_{\pm 0.003}$ $0.873_{\pm 0.077}$ $0.040_{\pm 0.000}$ $0.812_{\pm 0.073}$ $0.938_{\pm 0.027}$ $0.709_{\pm 0.377}$ $0.973_{\pm 0.006}$ $0.811_{\pm 0.031}$ $0.836_{\pm 0.108}$
boston $0.179_{\pm 0.084}$ $0.099_{\pm 0.071}$ $0.272_{\pm 0.192}$ $0.539_{\pm 0.242}$ $0.035_{\pm 0.000}$ $0.261_{\pm 0.075}$ $0.333_{\pm 0.377}$ $0.060_{\pm 0.031}$ $0.240_{\pm 0.251}$ $0.440_{\pm 0.019}$ $0.421_{\pm 0.039}$
breast $0.071_{\pm 0.060}$ $0.333_{\pm 0.468}$ $0.595_{\pm 0.415}$ $0.039_{\pm 0.026}$ $0.005_{\pm 0.000}$ $0.165_{\pm 0.193}$ $0.357_{\pm 0.256}$ $1.000_{\pm 0.000}$ $0.045_{\pm 0.002}$ $0.061_{\pm 0.019}$ $0.111_{\pm 0.071}$
cahouse $0.147_{\pm 0.038}$ $0.349_{\pm 0.246}$ $0.131_{\pm 0.185}$ $0.036_{\pm 0.005}$ $0.001_{\pm 0.000}$ $0.097_{\pm 0.073}$ $0.151_{\pm 0.120}$ $0.062_{\pm 0.035}$ $0.087_{\pm 0.058}$ $0.673_{\pm 0.209}$ $0.340_{\pm 0.157}$
credit $0.197_{\pm 0.068}$ $0.161_{\pm 0.124}$ $0.462_{\pm 0.131}$ $0.725_{\pm 0.261}$ $0.990_{\pm 0.000}$ $0.300_{\pm 0.120}$ $0.689_{\pm 0.246}$ $0.581_{\pm 0.194}$ $0.270_{\pm 0.077}$ $0.289_{\pm 0.121}$ $0.381_{\pm 0.043}$
diabetes $0.698_{\pm 0.230}$ $0.890_{\pm 0.036}$ $0.835_{\pm 0.187}$ $0.990_{\pm 0.005}$ $0.971_{\pm 0.000}$ $0.870_{\pm 0.147}$ $0.825_{\pm 0.093}$ $0.909_{\pm 0.049}$ $0.887_{\pm 0.078}$ $0.762_{\pm 0.044}$ $0.727_{\pm 0.036}$
iris $0.796_{\pm 0.127}$ $0.970_{\pm 0.039}$ $0.932_{\pm 0.032}$ $0.972_{\pm 0.038}$ $0.187_{\pm 0.000}$ $0.654_{\pm 0.353}$ $0.765_{\pm 0.149}$ $0.770_{\pm 0.118}$ $0.363_{\pm 0.176}$ $0.343_{\pm 0.124}$ $0.570_{\pm 0.118}$
qsar $0.668_{\pm 0.231}$ $0.546_{\pm 0.193}$ $0.669_{\pm 0.186}$ $0.485_{\pm 0.306}$ $0.063_{\pm 0.000}$ $0.750_{\pm 0.072}$ $0.664_{\pm 0.238}$ $0.351_{\pm 0.160}$ $0.794_{\pm 0.120}$ $0.951_{\pm 0.026}$ $0.952_{\pm 0.006}$
wdbc $0.821_{\pm 0.029}$ $0.647_{\pm 0.283}$ $0.909_{\pm 0.068}$ $1.000_{\pm 0.000}$ $0.591_{\pm 0.000}$ $0.784_{\pm 0.111}$ $0.853_{\pm 0.096}$ $0.307_{\pm 0.170}$ $0.911_{\pm 0.051}$ $0.977_{\pm 0.008}$ $0.956_{\pm 0.012}$
D.4 RDD Evaluation Results
Table 12: Performance comparison between different models in terms of RDD. AUC-ROC scores are reported, the smaller values mean that the synthetic records are harder to be detected. Best scores and second best scores are highlighted in bold with and without underscore, respectively. Equal values in the first 3 digits may be compared by the 4-th or 5-th digits.
Dataset ML ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
adult LN $0.794_{\pm 0.006}$ $0.632_{\pm 0.016}$ $0.946_{\pm 0.005}$ $\underline{\boldsymbol{0.585}}_{\pm 0.005}$ $0.714_{\pm 0.000}$ $\boldsymbol{0.602}_{\pm 0.000}$ $0.614_{\pm 0.031}$ $0.945_{\pm 0.003}$ $0.852_{\pm 0.001}$ $0.836_{\pm 0.001}$ $0.830_{\pm 0.001}$
RF $0.877_{\pm 0.003}$ $0.711_{\pm 0.010}$ $0.983_{\pm 0.007}$ $0.739_{\pm 0.002}$ $0.741_{\pm 0.001}$ $\underline{\boldsymbol{0.605}}_{\pm 0.002}$ $\boldsymbol{0.691}_{\pm 0.067}$ $0.978_{\pm 0.005}$ $0.871_{\pm 0.002}$ $0.868_{\pm 0.007}$ $0.858_{\pm 0.002}$
XGB $0.902_{\pm 0.003}$ $0.734_{\pm 0.011}$ $0.989_{\pm 0.004}$ $0.762_{\pm 0.002}$ $0.758_{\pm 0.001}$ $\underline{\boldsymbol{0.617}}_{\pm 0.002}$ $\boldsymbol{0.714}_{\pm 0.068}$ $0.985_{\pm 0.003}$ $0.894_{\pm 0.002}$ $0.874_{\pm 0.004}$ $0.859_{\pm 0.001}$
bank LN $0.736_{\pm 0.004}$ $0.607_{\pm 0.013}$ $0.818_{\pm 0.013}$ $\boldsymbol{0.581}_{\pm 0.004}$ $0.700_{\pm 0.000}$ $0.613_{\pm 0.001}$ $0.588_{\pm 0.025}$ $0.826_{\pm 0.021}$ $0.664_{\pm 0.017}$ $0.608_{\pm 0.010}$ $\underline{\boldsymbol{0.578}}_{\pm 0.006}$
RF $0.869_{\pm 0.014}$ $0.745_{\pm 0.007}$ $0.961_{\pm 0.006}$ $0.783_{\pm 0.003}$ $0.824_{\pm 0.000}$ $\underline{\boldsymbol{0.668}}_{\pm 0.002}$ $\boldsymbol{0.742}_{\pm 0.060}$ $0.968_{\pm 0.006}$ $0.995_{\pm 0.001}$ $0.991_{\pm 0.001}$ $0.990_{\pm 0.000}$
XGB $0.901_{\pm 0.008}$ $0.795_{\pm 0.008}$ $0.980_{\pm 0.002}$ $0.821_{\pm 0.007}$ $0.842_{\pm 0.001}$ $\underline{\boldsymbol{0.704}}_{\pm 0.002}$ $\boldsymbol{0.784}_{\pm 0.038}$ $0.983_{\pm 0.002}$ $0.997_{\pm 0.001}$ $0.992_{\pm 0.001}$ $0.990_{\pm 0.000}$
boston LN $0.579_{\pm 0.019}$ $0.741_{\pm 0.041}$ $0.830_{\pm 0.075}$ $0.994_{\pm 0.004}$ $0.674_{\pm 0.000}$ $0.583_{\pm 0.025}$ $0.625_{\pm 0.026}$ $0.614_{\pm 0.014}$ $0.547_{\pm 0.040}$ $\boldsymbol{0.578}_{\pm 0.006}$ $\underline{\boldsymbol{0.568}}_{\pm 0.007}$
RF $0.938_{\pm 0.014}$ $0.994_{\pm 0.002}$ $0.993_{\pm 0.003}$ $1.000_{\pm 0.000}$ $\boldsymbol{0.884}_{\pm 0.002}$ $\underline{\boldsymbol{0.651}}_{\pm 0.011}$ $0.963_{\pm 0.026}$ $0.993_{\pm 0.002}$ $0.995_{\pm 0.003}$ $0.986_{\pm 0.004}$ $0.987_{\pm 0.005}$
XGB $0.922_{\pm 0.017}$ $0.995_{\pm 0.001}$ $0.993_{\pm 0.005}$ $1.000_{\pm 0.000}$ $\boldsymbol{0.862}_{\pm 0.016}$ $\underline{\boldsymbol{0.642}}_{\pm 0.005}$ $0.978_{\pm 0.010}$ $0.996_{\pm 0.001}$ $0.996_{\pm 0.002}$ $0.993_{\pm 0.007}$ $0.995_{\pm 0.003}$
breast LN $0.605_{\pm 0.006}$ $0.723_{\pm 0.014}$ $0.690_{\pm 0.040}$ $0.619_{\pm 0.013}$ $0.685_{\pm 0.000}$ $0.658_{\pm 0.020}$ $0.695_{\pm 0.032}$ $0.702_{\pm 0.046}$ $0.615_{\pm 0.005}$ $\underline{\boldsymbol{0.573}}_{\pm 0.012}$ $\boldsymbol{0.582}_{\pm 0.020}$
RF $0.615_{\pm 0.014}$ $0.921_{\pm 0.041}$ $0.743_{\pm 0.021}$ $0.623_{\pm 0.006}$ $0.668_{\pm 0.001}$ $0.659_{\pm 0.013}$ $0.670_{\pm 0.019}$ $0.823_{\pm 0.012}$ $0.598_{\pm 0.006}$ $\underline{\boldsymbol{0.586}}_{\pm 0.011}$ $\boldsymbol{0.586}_{\pm 0.016}$
XGB $0.604_{\pm 0.011}$ $0.922_{\pm 0.041}$ $0.751_{\pm 0.026}$ $0.611_{\pm 0.005}$ $0.652_{\pm 0.017}$ $0.650_{\pm 0.024}$ $0.700_{\pm 0.018}$ $0.850_{\pm 0.017}$ $0.577_{\pm 0.005}$ $\underline{\boldsymbol{0.556}}_{\pm 0.009}$ $\boldsymbol{0.568}_{\pm 0.016}$
cahouse LN $\boldsymbol{0.517}_{\pm 0.003}$ $0.636_{\pm 0.063}$ $0.697_{\pm 0.017}$ $0.523_{\pm 0.005}$ $0.605_{\pm 0.000}$ $0.538_{\pm 0.002}$ $\underline{\boldsymbol{0.515}}_{\pm 0.004}$ $0.584_{\pm 0.017}$ $0.576_{\pm 0.010}$ $0.538_{\pm 0.002}$ $0.520_{\pm 0.005}$
RF $0.926_{\pm 0.003}$ $0.934_{\pm 0.007}$ $0.968_{\pm 0.011}$ $0.896_{\pm 0.003}$ $\boldsymbol{0.771}_{\pm 0.001}$ $\underline{\boldsymbol{0.658}}_{\pm 0.019}$ $0.949_{\pm 0.002}$ $0.940_{\pm 0.002}$ $0.936_{\pm 0.006}$ $0.875_{\pm 0.013}$ $0.816_{\pm 0.038}$
XGB $0.923_{\pm 0.004}$ $0.967_{\pm 0.007}$ $0.987_{\pm 0.003}$ $0.864_{\pm 0.001}$ $\boldsymbol{0.778}_{\pm 0.001}$ $\underline{\boldsymbol{0.672}}_{\pm 0.015}$ $0.891_{\pm 0.010}$ $0.969_{\pm 0.003}$ $0.985_{\pm 0.001}$ $0.934_{\pm 0.006}$ $0.903_{\pm 0.001}$
credit LN $0.505_{\pm 0.007}$ $0.634_{\pm 0.066}$ $0.654_{\pm 0.035}$ $0.663_{\pm 0.013}$ $0.708_{\pm 0.000}$ $0.639_{\pm 0.016}$ $0.812_{\pm 0.034}$ $0.688_{\pm 0.027}$ $0.658_{\pm 0.028}$ $\underline{\boldsymbol{0.485}}_{\pm 0.018}$ $\boldsymbol{0.491}_{\pm 0.002}$
RF $0.858_{\pm 0.003}$ $0.910_{\pm 0.074}$ $0.917_{\pm 0.017}$ $0.794_{\pm 0.026}$ $0.816_{\pm 0.001}$ $0.655_{\pm 0.016}$ $0.785_{\pm 0.038}$ $0.921_{\pm 0.011}$ $0.669_{\pm 0.023}$ $\underline{\boldsymbol{0.522}}_{\pm 0.015}$ $\boldsymbol{0.564}_{\pm 0.010}$
XGB $0.866_{\pm 0.004}$ $0.920_{\pm 0.083}$ $0.927_{\pm 0.013}$ $0.799_{\pm 0.021}$ $0.835_{\pm 0.005}$ $0.678_{\pm 0.010}$ $0.856_{\pm 0.017}$ $0.942_{\pm 0.008}$ $0.707_{\pm 0.036}$ $\boldsymbol{0.540}_{\pm 0.005}$ $\underline{\boldsymbol{0.526}}_{\pm 0.010}$
diabetes LN $\boldsymbol{0.547}_{\pm 0.009}$ $0.629_{\pm 0.045}$ $0.767_{\pm 0.026}$ $0.714_{\pm 0.012}$ $0.727_{\pm 0.000}$ $0.605_{\pm 0.009}$ $\underline{\boldsymbol{0.469}}_{\pm 0.021}$ $0.573_{\pm 0.018}$ $0.502_{\pm 0.036}$ $0.550_{\pm 0.012}$ $0.570_{\pm 0.014}$
RF $0.725_{\pm 0.021}$ $0.845_{\pm 0.016}$ $0.924_{\pm 0.038}$ $0.865_{\pm 0.017}$ $0.811_{\pm 0.001}$ $\boldsymbol{0.610}_{\pm 0.022}$ $\underline{\boldsymbol{0.603}}_{\pm 0.022}$ $0.953_{\pm 0.009}$ $0.743_{\pm 0.015}$ $0.713_{\pm 0.004}$ $0.753_{\pm 0.005}$
XGB $0.754_{\pm 0.018}$ $0.844_{\pm 0.014}$ $0.948_{\pm 0.019}$ $0.868_{\pm 0.011}$ $0.814_{\pm 0.003}$ $\underline{\boldsymbol{0.659}}_{\pm 0.027}$ $\boldsymbol{0.706}_{\pm 0.012}$ $0.977_{\pm 0.009}$ $0.722_{\pm 0.012}$ $0.722_{\pm 0.022}$ $0.738_{\pm 0.007}$
iris LN $0.518_{\pm 0.077}$ $0.796_{\pm 0.040}$ $0.867_{\pm 0.002}$ $0.806_{\pm 0.111}$ $0.534_{\pm 0.000}$ $0.619_{\pm 0.049}$ $0.485_{\pm 0.081}$ $0.510_{\pm 0.055}$ $0.494_{\pm 0.048}$ $\underline{\boldsymbol{0.433}}_{\pm 0.037}$ $\boldsymbol{0.473}_{\pm 0.050}$
RF $0.877_{\pm 0.023}$ $0.989_{\pm 0.008}$ $0.969_{\pm 0.007}$ $0.999_{\pm 0.002}$ $\boldsymbol{0.628}_{\pm 0.029}$ $\underline{\boldsymbol{0.549}}_{\pm 0.084}$ $0.896_{\pm 0.067}$ $0.954_{\pm 0.014}$ $0.988_{\pm 0.012}$ $0.996_{\pm 0.003}$ $0.969_{\pm 0.005}$
XGB $0.931_{\pm 0.037}$ $0.984_{\pm 0.010}$ $0.944_{\pm 0.003}$ $0.998_{\pm 0.003}$ $\underline{\boldsymbol{0.570}}_{\pm 0.009}$ $\boldsymbol{0.613}_{\pm 0.037}$ $0.896_{\pm 0.052}$ $0.975_{\pm 0.006}$ $0.992_{\pm 0.005}$ $0.997_{\pm 0.002}$ $0.981_{\pm 0.022}$
qsar LN $0.659_{\pm 0.025}$ $0.838_{\pm 0.031}$ $0.930_{\pm 0.013}$ $1.000_{\pm 0.000}$ $0.821_{\pm 0.000}$ $0.582_{\pm 0.022}$ $0.770_{\pm 0.016}$ $0.801_{\pm 0.034}$ $0.542_{\pm 0.013}$ $\underline{\boldsymbol{0.522}}_{\pm 0.004}$ $\boldsymbol{0.531}_{\pm 0.005}$
RF $0.972_{\pm 0.006}$ $1.000_{\pm 0.000}$ $0.999_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.979_{\pm 0.001}$ $\underline{\boldsymbol{0.677}}_{\pm 0.053}$ $\boldsymbol{0.946}_{\pm 0.018}$ $0.998_{\pm 0.001}$ $0.995_{\pm 0.001}$ $0.968_{\pm 0.011}$ $0.976_{\pm 0.007}$
XGB $0.984_{\pm 0.007}$ $1.000_{\pm 0.000}$ $0.999_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.989_{\pm 0.002}$ $\underline{\boldsymbol{0.761}}_{\pm 0.014}$ $\boldsymbol{0.975}_{\pm 0.003}$ $0.998_{\pm 0.001}$ $0.997_{\pm 0.001}$ $0.995_{\pm 0.002}$ $0.998_{\pm 0.000}$
wdbc LN $\boldsymbol{0.517}_{\pm 0.008}$ $0.832_{\pm 0.049}$ $0.961_{\pm 0.020}$ $0.994_{\pm 0.004}$ $0.675_{\pm 0.000}$ $0.564_{\pm 0.060}$ $0.658_{\pm 0.018}$ $0.743_{\pm 0.024}$ $0.550_{\pm 0.033}$ $\underline{\boldsymbol{0.511}}_{\pm 0.005}$ $0.521_{\pm 0.016}$
RF $0.948_{\pm 0.007}$ $1.000_{\pm 0.000}$ $0.993_{\pm 0.006}$ $1.000_{\pm 0.000}$ $0.965_{\pm 0.003}$ $0.695_{\pm 0.046}$ $0.616_{\pm 0.057}$ $0.989_{\pm 0.001}$ $0.653_{\pm 0.060}$ $\underline{\boldsymbol{0.536}}_{\pm 0.012}$ $\boldsymbol{0.543}_{\pm 0.020}$
XGB $0.939_{\pm 0.006}$ $0.999_{\pm 0.000}$ $0.990_{\pm 0.008}$ $1.000_{\pm 0.000}$ $0.968_{\pm 0.004}$ $0.709_{\pm 0.043}$ $0.677_{\pm 0.065}$ $0.994_{\pm 0.001}$ $0.650_{\pm 0.011}$ $\underline{\boldsymbol{0.500}}_{\pm 0.003}$ $\boldsymbol{0.502}_{\pm 0.003}$
Avg. $0.777_{\pm 0.166}$ $0.843_{\pm 0.136}$ $0.904_{\pm 0.107}$ $0.830_{\pm 0.159}$ $0.767_{\pm 0.120}$ $\underline{\boldsymbol{0.638}}_{\pm 0.050}$ $0.742_{\pm 0.149}$ $0.872_{\pm 0.150}$ $0.765_{\pm 0.186}$ $0.726_{\pm 0.207}$ $\boldsymbol{0.726}_{\pm 0.200}$
Table 12 shows the result for RDD. Since TabTreeFormernaïvely use mean value as the reverse quantization step, it is expected that as long as a detector is able to detect these specific recovered quantized values, TabTreeFormer-generated records can be detected. Even so, TabTreeFormer still outperforms all baselines except for REaLTabFormer (Solatorio & Dupriez, 2023).
D.5 Computation Time
Table 13: Comparison between different models in terms of time taken for training and generation. The values are in unit of seconds.
Dataset Action ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
adult Train $258.758_{\pm 55.030}$ $787.379_{\pm 3.334}$ $753.259_{\pm 104.080}$ $285.153_{\pm 0.410}$ $10397.040_{\pm 9.203}$ $1875.904_{\pm 34.663}$ $1199.823_{\pm 146.744}$ $1111.927_{\pm 2.790}$ $166.807_{\pm 31.955}$ $235.554_{\pm 71.542}$ $606.790_{\pm 277.828}$
Generate $41.029_{\pm 0.363}$ $0.429_{\pm 0.001}$ $0.112_{\pm 0.001}$ $63.388_{\pm 0.083}$ $138.023_{\pm 0.205}$ $89.466_{\pm 4.301}$ $2.409_{\pm 0.010}$ $0.357_{\pm 0.005}$ $30.822_{\pm 3.553}$ $35.275_{\pm 6.059}$ $96.027_{\pm 37.920}$
bank Train $121.480_{\pm 39.164}$ $737.314_{\pm 0.138}$ $544.578_{\pm 62.697}$ $252.016_{\pm 0.266}$ $10295.769_{\pm 21.720}$ $835.167_{\pm 1.619}$ $1865.473_{\pm 418.665}$ $878.170_{\pm 122.998}$ $228.772_{\pm 62.554}$ $225.707_{\pm 39.181}$ $744.726_{\pm 221.487}$
Generate $42.208_{\pm 5.408}$ $1.126_{\pm 0.006}$ $0.733_{\pm 0.001}$ $55.038_{\pm 0.077}$ $149.872_{\pm 0.104}$ $106.036_{\pm 0.080}$ $3.009_{\pm 0.063}$ $0.975_{\pm 0.003}$ $36.731_{\pm 3.457}$ $37.163_{\pm 3.021}$ $117.167_{\pm 30.088}$
boston Train $1.184_{\pm 0.368}$ $8.396_{\pm 0.048}$ $14.945_{\pm 5.245}$ $4.499_{\pm 0.047}$ $121.619_{\pm 0.090}$ $49.693_{\pm 3.175}$ $862.625_{\pm 48.241}$ $7.158_{\pm 0.387}$ $94.498_{\pm 42.836}$ $224.217_{\pm 30.311}$ $432.354_{\pm 180.912}$
Generate $1.184_{\pm 0.269}$ $0.083_{\pm 0.001}$ $0.062_{\pm 0.000}$ $2.332_{\pm 0.459}$ $2.422_{\pm 0.003}$ $3.059_{\pm 0.003}$ $0.156_{\pm 0.003}$ $0.066_{\pm 0.000}$ $0.372_{\pm 0.094}$ $0.694_{\pm 0.055}$ $1.275_{\pm 0.379}$
breast Train $0.585_{\pm 0.015}$ $8.479_{\pm 0.026}$ $13.571_{\pm 2.035}$ $6.136_{\pm 0.099}$ $153.352_{\pm 0.134}$ $25.537_{\pm 1.419}$ $1037.126_{\pm 281.166}$ $6.340_{\pm 0.925}$ $64.589_{\pm 26.284}$ $207.280_{\pm 0.096}$ $537.723_{\pm 206.786}$
Generate $0.876_{\pm 0.013}$ $0.056_{\pm 0.008}$ $0.051_{\pm 0.000}$ $3.123_{\pm 0.007}$ $3.434_{\pm 0.016}$ $0.759_{\pm 0.003}$ $0.249_{\pm 0.003}$ $0.055_{\pm 0.000}$ $0.315_{\pm 0.050}$ $0.527_{\pm 0.002}$ $1.315_{\pm 0.403}$
cahouse Train $109.471_{\pm 13.706}$ $280.228_{\pm 0.142}$ $311.628_{\pm 21.038}$ $96.491_{\pm 0.350}$ $4373.950_{\pm 56.089}$ $596.523_{\pm 51.206}$ $1838.247_{\pm 200.652}$ $208.611_{\pm 25.684}$ $267.307_{\pm 29.842}$ $340.992_{\pm 76.571}$ $946.903_{\pm 102.584}$
Generate $8.087_{\pm 0.220}$ $0.839_{\pm 0.016}$ $0.690_{\pm 0.011}$ $16.056_{\pm 0.012}$ $57.584_{\pm 0.061}$ $69.522_{\pm 0.044}$ $2.676_{\pm 0.008}$ $0.787_{\pm 0.006}$ $7.774_{\pm 0.457}$ $9.725_{\pm 1.540}$ $29.610_{\pm 2.509}$
credit Train $2.262_{\pm 0.089}$ $8.400_{\pm 0.035}$ $32.463_{\pm 8.548}$ $8.102_{\pm 0.215}$ $277.555_{\pm 0.388}$ $63.648_{\pm 0.758}$ $726.006_{\pm 67.662}$ $13.246_{\pm 1.809}$ $152.310_{\pm 23.655}$ $250.729_{\pm 75.271}$ $446.486_{\pm 161.523}$
Generate $2.198_{\pm 0.088}$ $0.043_{\pm 0.000}$ $0.035_{\pm 0.000}$ $4.526_{\pm 0.007}$ $5.845_{\pm 0.008}$ $2.407_{\pm 0.171}$ $0.167_{\pm 0.003}$ $0.041_{\pm 0.000}$ $0.541_{\pm 0.153}$ $1.062_{\pm 0.179}$ $2.018_{\pm 0.586}$
diabetes Train $1.100_{\pm 0.094}$ $8.524_{\pm 0.407}$ $20.414_{\pm 2.831}$ $5.799_{\pm 1.191}$ $130.793_{\pm 0.077}$ $37.372_{\pm 3.118}$ $758.140_{\pm 41.333}$ $7.944_{\pm 1.272}$ $69.127_{\pm 35.422}$ $74.799_{\pm 18.217}$ $560.416_{\pm 149.115}$
Generate $0.943_{\pm 0.047}$ $0.047_{\pm 0.001}$ $0.036_{\pm 0.000}$ $2.329_{\pm 0.822}$ $1.496_{\pm 0.008}$ $1.504_{\pm 0.007}$ $0.139_{\pm 0.000}$ $0.039_{\pm 0.000}$ $0.412_{\pm 0.044}$ $0.427_{\pm 0.023}$ $1.580_{\pm 0.318}$
iris Train $0.153_{\pm 0.002}$ $7.267_{\pm 0.021}$ $6.538_{\pm 0.504}$ $3.485_{\pm 0.210}$ $24.975_{\pm 0.018}$ $15.369_{\pm 1.242}$ $620.864_{\pm 35.603}$ $2.811_{\pm 0.518}$ $225.728_{\pm 194.228}$ $406.082_{\pm 0.376}$ $482.266_{\pm 0.176}$
Generate $0.690_{\pm 0.010}$ $0.033_{\pm 0.006}$ $0.016_{\pm 0.000}$ $6.208_{\pm 2.124}$ $0.360_{\pm 0.015}$ $0.225_{\pm 0.001}$ $0.098_{\pm 0.002}$ $0.016_{\pm 0.001}$ $0.124_{\pm 0.054}$ $0.198_{\pm 0.002}$ $0.331_{\pm 0.010}$
qsar Train $18.992_{\pm 1.985}$ $18.389_{\pm 0.217}$ $64.331_{\pm 11.582}$ $7.369_{\pm 0.156}$ $566.658_{\pm 23.451}$ $152.950_{\pm 0.819}$ $694.890_{\pm 76.489}$ $23.013_{\pm 2.183}$ $132.605_{\pm 45.722}$ $297.779_{\pm 59.962}$ $739.373_{\pm 278.301}$
Generate $2.842_{\pm 0.012}$ $0.206_{\pm 0.006}$ $0.192_{\pm 0.000}$ $2.825_{\pm 0.826}$ $46.712_{\pm 3.062}$ $10.997_{\pm 0.072}$ $0.336_{\pm 0.005}$ $0.203_{\pm 0.001}$ $1.059_{\pm 0.032}$ $2.375_{\pm 0.210}$ $5.404_{\pm 1.685}$
wdbc Train $5.344_{\pm 0.860}$ $10.795_{\pm 0.089}$ $24.175_{\pm 7.507}$ $4.249_{\pm 0.056}$ $265.340_{\pm 2.411}$ $183.908_{\pm 1.047}$ $631.546_{\pm 68.070}$ $9.356_{\pm 1.236}$ $132.480_{\pm 48.560}$ $235.600_{\pm 110.062}$ $504.627_{\pm 29.714}$
Generate $1.410_{\pm 0.007}$ $0.159_{\pm 0.001}$ $0.154_{\pm 0.002}$ $1.934_{\pm 0.010}$ $65.821_{\pm 1.106}$ $9.936_{\pm 0.017}$ $0.276_{\pm 0.005}$ $0.159_{\pm 0.000}$ $0.606_{\pm 0.019}$ $1.252_{\pm 0.180}$ $2.231_{\pm 0.084}$
Avg. Train $51.933_{\pm 86.361}$ $187.517_{\pm 314.694}$ $178.590_{\pm 268.489}$ $67.330_{\pm 110.073}$ $2660.705_{\pm 4255.243}$ $383.607_{\pm 593.187}$ $1023.474_{\pm 472.880}$ $226.858_{\pm 413.306}$ $153.422_{\pm 69.440}$ $249.874_{\pm 87.633}$ $600.166_{\pm 163.582}$
Generate $10.147_{\pm 16.730}$ $0.302_{\pm 0.384}$ $0.208_{\pm 0.271}$ $15.776_{\pm 23.351}$ $47.157_{\pm 56.920}$ $29.391_{\pm 41.741}$ $0.952_{\pm 1.216}$ $0.270_{\pm 0.341}$ $7.876_{\pm 13.911}$ $8.870_{\pm 14.692}$ $25.696_{\pm 43.802}$
TabTreeFormer is trained by limiting the total number of training steps instead of the number of epochs, so the training time of it on smaller datasets are usually longer.
D.6 Raw Results for Ablation Study
Table 14: Ablation studies.
Dataset credit diabetes
ML LN RF XGB LN RF XGB
TTF $0.823_{\pm 0.022}$ $0.811_{\pm 0.014}$ $0.816_{\pm 0.015}$ $0.889_{\pm 0.002}$ $0.865_{\pm 0.009}$ $0.863_{\pm 0.006}$
w/o TBM $0.761_{\pm 0.030}$ $0.747_{\pm 0.023}$ $0.726_{\pm 0.047}$ $0.880_{\pm 0.002}$ $0.848_{\pm 0.004}$ $0.824_{\pm 0.010}$
w/o DQT $0.725_{\pm 0.013}$ $0.807_{\pm 0.002}$ $0.813_{\pm 0.010}$ $0.881_{\pm 0.002}$ $0.863_{\pm 0.007}$ $0.861_{\pm 0.002}$
w/o OCEL $0.839_{\pm 0.009}$ $0.813_{\pm 0.024}$ $0.815_{\pm 0.017}$ $0.881_{\pm 0.002}$ $0.847_{\pm 0.006}$ $0.839_{\pm 0.007}$
Table 14 shows the raw results for ablation study.
Appendix E Comparison to Baselines without Public Code
Table 15: Average MLE relative error of the synthetic data compared to the real data. “Ours” refer to the results from this paper. The other row names are the source paper of the values reported in the row.
ARF CTAB+ CTGAN TabDDPM GReaT RTF TabSyn TVAE TabMT TabDiff TTF-S TTF-M TTF-L
Ours $0.057_{\pm 0.077}$ $0.244_{\pm 0.274}$ $0.099_{\pm 0.137}$ $0.483_{\pm 1.214}$ $0.110_{\pm 0.170}$ $0.027_{\pm 0.040}$ $0.027_{\pm 0.029}$ $0.095_{\pm 0.142}$ $0.029_{\pm 0.040}$ $0.017_{\pm 0.023}$ $0.016_{\pm 0.023}$
RTF - $0.181_{\pm 0.187}$ $0.062_{\pm 0.060}$ $0.118_{\pm 0.151}$ $0.050_{\pm 0.038}$ $0.130_{\pm 0.114}$
TabMT - $0.167_{\pm 0.167}$ $0.031_{\pm 0.041}$ $0.054_{\pm 0.034}$ $0.135_{\pm 0.126}$ $0.021_{\pm 0.021}$
TabDiff - - $0.237_{\pm 0.398}$ $0.784_{\pm 1.765}$ $0.133_{\pm 0.230}$ $0.068_{\pm 0.134}$ $0.201_{\pm 0.279}$ $0.058_{\pm 0.112}$
There are two recent tabular data generation works, TabMT (Gulati & Roysdon, 2024) and TabDiff (Shi et al., 2024), that have not published their code by the time of the composition of this paper. Unfortunately, although all papers report MLE scores, which can be a basis of comparison, the values from these papers cannot be directly taken for comparison because of the different settings of evaluation, including datasets, splits, ML model, and evaluation metric. However, fortunately, these papers report the performance of a few other common baselines and report the performance of the real data too. Thus, we use the reported values in these papers as indirect evidence for comparing these models’ performances. The values are shown in Table 15. While we acknowledge that experiments are the most solid way to showcase superior performance, this section mainly serves as a temporary qualitative analysis when these works’ code is not public yet.
TabMT. TabTreeFormer wins over CTAB-GAN+ and TabDPPM by a much greater margin [^2] By “margin”, we compare the change ratio instead of absolute relative error values, as different papers tend to have values reported at different scales and variance (compared between different models)., wins over TVAE by a similar margin, and wins over REaLTabFormer by a lower margin. Moreover, the results for REaLTabFormer reported by TabMT are not done by reproduced experiments, but directly taken from REaLTabFormer paper. Experiments done separately, possibly on different splits, may result in an non-negligible change in the MLE scores, which is also acknowledged by TabMT authors. In addition, TabMT’s result is reported on 15 datasets while REaLTabFormer’s result is reported on 6 datasets only, and the 6 datasets reported for REaLTabFormer are harder than the rest 9 datasets by a higher MLE relative error over all the other reported baselines in that paper. Therefore, TabTreeFormer is likely to outperform TabMT.
TabDiff. TabTreeFormer wins over all the reported baselines with a greater margin. Thus, it is also vrey likely that TabTreeFormer outperforms TabDiff.
Appendix F Limitations and Future Work
Optimizing tree-based model. TabTreeFormer uses LightGBM fitted on the target column to introduce tabular inductive biases from tree-based models, While this is effective and easy to implement, modified tree-based models catered for guiding the generation may further improve the performance, such as using trees from a generative tree-based model (e.g. ARF), and selecting a better target column using some heuristics.
Optimizing de-quantization of tokenization. The dual-quantization tokenization may be further optimized with some sampling around the quantile values if use case requires smoother values in continuous values (e.g. for RDD evaluation, recall Appendix D.4).
Masked vs. Auto-regressive transformers. TabMT has demonstrated the effectiveness of a masked transformer with iterative decoding for tabular data generation, but TabMT does not provide ablation study result on masked vs. auto-regressive transformers. Thus, this approach may be used in place of a causal language model to further improve synthetic data quality. Further exploration could be useful.
Extension to general tabular models. The idea of introducing tree-based models to transformers on tabular data may not be limited to generation task, but also applies to tasks like classification and regression. Future works may explore the extendiblity of this idea to other tasks.

### v5

@@
 Meanwhile, they suffer from poor scalability and efficiency due to quadratic computational complexity.
-In this paper, we propose TabTreeFormer, a hybrid transformer architecture that incorporates a tree-based model that retains tabular-specific inductive biases of non-smooth and potentially low-correlated patterns due to its discreteness and non-rotational invariance, and hence enhances the fidelity and utility of synthetic data.
-In addition, we devise a dual-quantization tokenizer to capture the multimodal continuous distribution and further facilitate the learning of numerical value distribution.
-Moreover, our proposed tokenizer reduces the vocabulary size and sequence length due to the limited dimension-wise semantic meaning and training set size of tabular data, rendering a significant model size shrink without sacrificing the capability of the transformer model.
-We evaluate TabTreeFormer on 10 datasets against multiple generative models on various metrics; our experimental results show that TabTreeFormer achieves superior fidelity, utility, privacy, and efficiency.
+In this paper, we propose TabTreeFormer, a hybrid transformer architecture that incorporates inductive biases of tree-based models (i.e., non-smoothness and non-rotational invariance) to process the potentially discrete and low-correlated patterns in tabular data, and hence enhances the fidelity and utility of synthetic data.
+In addition, we devise a novel tokenizer to capture the multimodal continuous distribution and further facilitate the learning of numerical value distribution.
+Moreover, our proposed tokenizer reduces the vocabulary size and sequence length based on the complexity of tabular data, shrinking the model size significantly without sacrificing the capability of the transformer model.
+We evaluate TabTreeFormer on 10 datasets against multiple generative models in various metrics; our experimental results show that TabTreeFormer achieves superior fidelity, utility, privacy, and efficiency.
 Our best model yields a 40% utility improvement with 1/16 of the baseline model size.
+Code is available at: https://anonymous.4open.science/r/tabtreeformer-9585.
 Tabular Data, Generative Models
@@
 Figure 1: Performance comparison of TabTreeFormer (Ours) and other representative tabular generative models in terms of fidelity, utility (i.e., MLE), privacy, and efficiency.
-TabTreeFormer achieves the best overall performance and outperforms other models on almost all sole metrics.
-Tables are one of the most prevalent data modalities in real-world applications (healthcare (Dash et al., 2019), financial services (Assefa et al., 2021), etc.), yet are heavily under-exploited due to privacy concerns (EU, 2016).
+TabTreeFormer outperforms other models on almost all metrics.
+Tables are one of the most prevalent data modalities in real-world applications (e.g., healthcare (Dash et al., 2019), financial services (Assefa et al., 2021), etc.), yet are heavily under-exploited due to privacy concerns (EU, 2016).
 Fortunately, synthetic data offers an alternative to the utilization of tabular data by modeling the characteristics of real data and reducing the risk of data breach (PDPC, 2024), thus drawing considerable attention in recent years.
-The state-of-the-art research shows that transformers such as autoregressive transformers (Borisov et al., 2023), masked transformers (Gulati & Roysdon, 2024), and diffusion models with transformers (Zhang et al., 2024) have achieved impressive fidelity, utility, and privacy in tabular data generation, enabling synthetic data to empower a variety of fields (Hernandez et al., 2022; Assefa et al., 2021).
+State-of-the-art research shows that transformers such as autoregressive transformers (Borisov et al., 2023), masked transformers (Gulati & Roysdon, 2023), and diffusion models with transformers (Zhang et al., 2024) have achieved impressive fidelity, utility, and privacy in tabular data generation, allowing synthetic data to empower a variety of fields (Hernandez et al., 2022; Assefa et al., 2021).
 However, unlike the application of transformers in other research areas such as computer vision (Dosovitskiy et al., 2021) and natural language processing (Vaswani et al., 2017), existing transformer models for synthetic tabular data often overlook domain-specific priors (i.e., inductive biases).
@@
   1. 1.
-    Are tabular-specific inductive biases beneficial to the performance of tabular generative models?
+    What inductive biases are beneficial to the fidelity and utility of synthetically generated tabular data?
   2. 2.
-    Can one exploit unique characteristics of tabular data to improve model scalability and efficiency?
-To answer the questions, we propose TabTreeFormer, a generative hybrid transformer that incorporates a tree-based model and a deliberately devised tokenizer to process and introduce tabular-specific inductive biases.
-We also leverage the limited per-dimension semantic meaning and limited training set size of typical tabular data to enable drastically reduced model size by our tokenizer design.
-Inspired by the superior performance of tree-based models on tabular tasks (Shwartz-Ziv & Armon, 2022; Gorishniy et al., 2024), we propose to employ them on capturing tabular-specific inductive biases of non-smoothness and potentially low-correlation (Grinsztajn et al., 2024).
-Non-smoothness is attributed to the existence of discrete features and non-smooth relations between discrete and/or continuous features.
-Neural networks are biased to overly smooth solutions while tree-based models are exceptionally good at capturing the non-smooth patterns since they learn piece-wise constant functions rather than low-frequency functions (i.e., the functions neural networks tend to bias toward) (Rahaman et al., 2019).
-Low-correlated features includes uninformative features that contribute trivially to downstream tasks such as classification and regression.
-Due to the non-rotationally invariant nature of tree-based models (Ng, 2004), trees are more robust against low-correlated features, while neural networks are biased towards the universal presence of correlation.
-Nevertheless, trees do not capture multimodal distribution in continuous features, which is a tabular inductive bias particularly essential for generative tasks (Xu et al., 2019; Shi et al., 2024).
-We then propose dual-quantization tokenization of numeric values, where the first quantization is a K-Means (MacQueen, 1967) quantization that captures multimodal distribution, and the second quantization is a quantile quantization that is used to present the precise value.
-Considering the structured nature of tabular data and limited semantic meaning in each dimension, we further optimize the tokenizer so that the vocabulary size and sequence length can be reduced significantly.
-Together with the limited of training data compared to vision or language tasks, this leads to room for reducing other hyperparameters such as hidden dimensions, and eventually the entire model size.
+    How can one exploit unique characteristics of tabular data to improve model scalability and efficiency?
+To answer the questions, we propose TabTreeFormer, a generative hybrid transformer incorporating a tree-based model and a deliberately devised tokenizer to process and introduce tabular-specific inductive biases.
+We also leverage the limited per-dimension semantic meaning (i.e., each dimension corresponds to at most one feature) of typical tabular data to enable a significant reduction in model size by our tokenizer design.
+Tree-based models excel at tabular classification and regression tasks (Shwartz-Ziv & Armon, 2022; Gorishniy et al., 2021), which is attributed to their inductive biases to capture tabular characteristics such as non-smoothness and potentially low-correlation (Grinsztajn et al., 2022).
+Non-smoothness is due to the existence of discrete features and non-smooth relations between discrete and/or continuous features.
+Tree-based models learn piece-wise constant functions rather than low-frequency functions (i.e., the functions neural networks tend to learn) (Rahaman et al., 2019), so that they are exceptionally good at capturing such non-smooth patterns.
+textitLow-correlated ifeatures nclude uninformative features that contribute trivially to downstream tasks such as classification and regression.
+Due to the non-rotationally invariant nature of tree-based models (Ng, 2004), trees are more robust against low-correlated features, while neural networks are biased towards stronger correlations.
+Inspired by the success of tree models on tabular tasks due to their inductive bias, we propose to employ such priors to facilitate the performance of tabular generative tasks.
+However, tree models are not panaceas as they do not inherently capture multimodal distribution in continuous features (i.e., features with probability density functions possessing multiple modes/peaks), which is an essential challenge for tabular generative tasks.
+Inspired by the prior work (Xu et al., 2019) that tackles the challenge using multimodal decomposition, we propose dual-quantization tokenization for numeric values, a variant of multimodal decomposition catered for transformers where the first quantization is a K-Means (MacQueen, 1967) quantization that captures the multimodal distribution, and the second quantization is a quantile quantization for more precise representation of numerical values.
+Moreover, we further optimize the tokenizer to significantly reduce vocabulary size and sequence length by considering the structured nature of tabular data and limited semantic meaning in each dimension (i.e., sole feature for each dimension).
+In addition, the limited complexity of training data in tables (compared to vision and language tasks) leads to room for reducing model complexity (e.g., hidden dimensions, number of heads, number of layers, etc.) and eventually slimming down the entire model size.
 Our contributions are summarized as follows:
-  • To the best of our knowledge, we are the first work that introduces tabular-specific inductive biases to improve the performance of tabular generative models.
-We propose a novel hybrid TFM coupled with a TBM to capture such inductive biases.
-  • We devise a dual-quantization tokenizer that models the multimodal distribution of continuous values in real data.
-This practice facilitates the learning of the numerical distribution and further improves the quality of synthetic tabular data.
+  • To the best of our knowledge, we are the first work that introduces tabular-specific inductive biases to transformers to improve tabular generative models.
+We propose a novel hybrid transformer coupled with a tree-based model to bring in such inductive biases.
+  • We devise a dual-quantization tokenizer that models the multimodal distribution of continuous values.
+This practice facilitates learning of numerical distribution and further improves the quality of synthetic data.
   • Employing the characteristics of tabular data, our tokenizer design also leads to a much smaller vocabulary size and sequence length, achieving a phenomenal model size reduction and efficiency enhancement.
-  • We evaluate TabTreeFormer across 10 datasets against multiple representative tabular generative models and show that TabTreeFormer outperforms other models in fidelity, utility, privacy, and efficiency, as shown in Fig.
+  • We evaluate TabTreeFormer across 10 datasets against 8 representative tabular generative models and show that TabTreeFormer outperforms other models in fidelity, utility, privacy, and efficiency, as shown in Fig.
 1.
@@
 Early works on tabular data generation predominantly utilized MLPs and CNNs as backbone architectures and GAN and VAE as generation methods (Park et al., 2018; Xu et al., 2019; Zhao et al., 2021, 2024).
-Recent works evolved to use transformers with auto-regression, masked modeling, and diffusion as generative paradigms (Borisov et al., 2023; Zhang et al., 2024; Gulati & Roysdon, 2024).
-For example, TabMT (Gulati & Roysdon, 2024) employs a masked transformer with ordered embedding and achieves outstanding utility in downstream tasks with improved scalability, and TabSyn (Zhang et al., 2024) encodes tabular data into a latent space and generates tabular data using diffusion model, demonstrating impressive fidelity and utility.
-Despite their enhanced performance, these works overlook the essential inductive biases such as non-smoothness and low-correlated features that are crucial to preserving the intrinsic characteristics of tabular data.
+Recent works evolved to use transformers with auto-regression, masked modeling, and diffusion as generative paradigms (Borisov et al., 2023; Zhang et al., 2024; Gulati & Roysdon, 2023).
+For example, TabMT (Gulati & Roysdon, 2023) employs a masked transformer with ordered embedding and achieves outstanding utility in downstream tasks with improved scalability, and TabSyn (Zhang et al., 2024) encodes tabular data into a latent space and generates tabular data using diffusion model, demonstrating impressive fidelity and utility.
+Despite their enhanced performance, these works overlook the essential inductive biases to capture non-smoothness and low-correlated features that are crucial to preserving the intrinsic characteristics of tabular data.
 Xu et al.
-(2019) finds that multimodal distribution in continuous features is a critical inductive bias for table generation.
+(2019) finds that multimodal distribution in continuous features is a critical aspect for tabular generative models’ inductive bias to capture.
 CTGAN (Xu et al., 2019) uses variational Gaussian mixture model (Bishop, 2006) to decompose multimodal values while TabDiff (Shi et al., 2024) introduces a multimodal stochastic sampler.
-In this work, we introduce multimodal distribution, into an auto-regressive transformer, which was absent in prior works.
+In this work, we devise a tokenizer that models the multimodal distribution and introduces the prior into the auto-regressive transformer.
 2.2 Tree-based Models for Tabular Data
-In recent years, tree-based gradient boosting algorithms have emerged as the dominant approach for tabular data tasks, particularly with the introduction of models like XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018).
-These tree-based models are ensembles of small and intuitive decision trees that naturally understand feature relations and interactions, thus achieving high predictive performance, efficiency, and scalability, making them the preferred choice in both academia and industry.
-While some recent research has shown promise of transformers in tabular data (Huang et al., 2020; Arik & Pfister, 2021; Hollmann et al., 2023) in certain scenarios, their computational cost and suboptimal performance make them still less favorable compared to tree-based models (Shwartz-Ziv & Armon, 2022; Gorishniy et al., 2024), which is attributed to tabular-specific inductive biases, including non-smoothness and potentially low-correlation, captured by tree-based models (Grinsztajn et al., 2024).
-In this work, we leverage these features of tree-based models and retain tabular-specific inductive bias to facilitate the training of our transformer model.
+Tree-based gradient boosting algorithms are the dominant approach for tabular tasks, notable examples are XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018).
+These tree-based models are ensembles of small and intuitive decision trees that naturally understand feature relations and interactions, thus achieving high predictive performance, efficiency, and scalability.
+Although recent research has shown promise of transformers in tabular data-related tasks (Huang et al., 2020; Arik & Pfister, 2021; Hollmann et al., 2023), their suboptimal performance caused by the absence of tabular-specific inductive biases makes them still less favorable to tree-based models in many scenarios (Shwartz-Ziv & Armon, 2022; Gorishniy et al., 2021).
+Meanwhile, tree-based generative models (Watson et al., 2023; McCarter, 2024), while being efficient, still fall short of synthetic data quality compared to deep neural networks like transformers.
+In this work, we take advantage of both models and employ tree-based models to introduce tabular-specific inductive bias and facilitate the learning of the transformer model.
 2.3 Tokenizers for Tabular Data
@@
 Borisov et al.
-(2023); Solatorio & Dupriez (2023); Gulati & Roysdon (2024) translate tabular data to a sequence of tokens (i.e., discrete space) by a tabular-specific tokenization method, and retain the embedding layer with moderate modifications, which is akin to natural language and more suitable for auto-regressive and masked transformers.
-In this paper, we opt for the second choice, to optimize the tokenizer for tabular data.
-3 TabTreeFormer
-Figure 2: Overview of TabTreeFormer.
-Tokens representing tree leaves are prepended to each row’s tokenized representation, which is also heavily optimized.
-Tree leaves and optimized data tokens introduce essential inductive biases to the tabular data generator.
-Recall that our goal is to improve the performance and efficiency of the generative model by introducing domain-specific inductive biases and shrinking the model size.
-Formally, let $\mathbf{X}$ be the tabular training dataset for a generative model with $n$ rows, $m_{d}$ discrete and $m_{c}$ continuous features, we aim to generate synthetic tabular data $\mathbf{X}^{\prime}$ similar to $\mathbf{X}$.
-To achieve this, we propose TabTreeFormer that consists of three building blocks: a tree-based model, a tokenizer, and a transformer model, as shown in Fig.
-2.
-The tree-based model captures some tabular-specific inductive biases, and the tokenizer is used to grasp the multimodal distribution while reducing vocabulary size and sequence length simultaneously.
-The transformer model learns the prior extracted from the tree-based model and tokenizer and produces high-quality synthetic data.
-Concrete accounts of training and generation algorithms are described in Algorithm 1 and 2.
+(2023); Solatorio & Dupriez (2023); Gulati & Roysdon (2023) translate tabular data to a sequence of tokens (i.e., discrete space) by a tabular-specific tokenization method, and retain the embedding layer with moderate modifications, which is akin to natural language and more suitable for transformers.
+In this paper, we opt for the second choice to optimize the tokenizer for tabular data.
 Algorithm 1 Training of TabTreeFormer
 1:  Input: Tabular dataset $\mathbf{X}$ ($n$ rows, $m_{d}$ discrete features, $m_{c}$ continuous features), $s$ total training steps
-2:  Output: Fitted tree-based model $\mathcal{T}$ and the leaf index matrix $\mathbf{J}$, data tokenizer $\mathcal{D}$, and transformer-based language model $\mathcal{G}$
+2:  Output: Fitted tree-based model $\mathcal{T}$ and the leaf index matrix $\mathbf{J}$, data tokenizer $\mathcal{D}$, and transformer $\mathcal{G}$
 […]
@@
 13:  end for
+3 TabTreeFormer
+Figure 2: Overview of TabTreeFormer (data flow from bottom to top).
+TabTreeFormer consists of 3 components: i) a tree-based model that introduces tabular-specific inductive biases; ii) a tokenizer that efficiently and compactly represents data while capturing multimodal distributions; iii) a transformer model that learns the priors extracted from the tree and tokenizer to generate high-quality synthetic data.
+Recall that our goal is to improve the performance and efficiency of the generative model by introducing domain-specific inductive biases and shrinking the model size.
+Formally, let $\mathbf{X}$ be the tabular training dataset for a generative model with $n$ rows, $m_{d}$ discrete and $m_{c}$ continuous features, we aim to generate synthetic tabular data $\mathbf{X}^{\prime}$ similar to $\mathbf{X}$.
+To achieve this, we propose TabTreeFormer that consists of three building blocks: a tree-based model, a tokenizer, and a transformer model, as shown in Fig.
+2.
+The tree-based model captures some tabular-specific inductive biases, and the tokenizer is used to grasp the multimodal distribution while reducing vocabulary size and sequence length simultaneously.
+The transformer model learns the priors extracted from the tree-based model and tokenizer to produce high-quality synthetic data.
+Concrete accounts of training and generation algorithms are described in Algorithm 1 and 2.
 Algorithm 2 Tabular data generation using TabTreeFormer
-1:  Input: Leaf index matrix $\mathbf{J}$, data tokenizer $\mathcal{D}$, transformer-based language model $\mathcal{G}$, and number of rows to be generated $n^{\prime}$
+1:  Input: Leaf index matrix $\mathbf{J}$, data tokenizer $\mathcal{D}$, transformer-based language model $\mathcal{G}$, and number of rows generate $n^{\prime}$
 2:  Output: Synthetic tabular dataset $\mathbf{X}^{\prime}$
@@
 2).
-This is achieved by augmenting each token sequence representing one table row with additional information derived from the leaf indices of a tree-based model.
+This is achieved by augmenting each token sequence representing one table row with additional information derived from the leaf indices of a tree-based model fitted on the prediction task defined on the dataset or a random feature if such a task is unavailable.
 Formally, let the fitted tree-based model be $\mathcal{T}$ with $T$ trees.
 Let $l_{i}\in\mathbb{N}$ be the number of leaves in $i$-th tree, where $i\in\{1,2,\dots,T\}$.
-On $\mathcal{T}$, we have the leaf index matrix $\mathbf{J}=[j_{ik}]\in\mathbb{N}^{n\times T}$ of $\mathbf{X}$ where $j_{ik}\in\{1,2,\dots,l_{k}\}$ represents the leaf index of $i$-th row $\mathbf{X}_{i}$ in the $k$-th tree.
-Leaf indices provide flattened positional information of each row in a tree, and hence inherently introduce tabular inductive biases captured by trees – non-smoothness and potentially low-correlations.
-With leaf indices prepended to actual data values, although the subsequent step is still a neural network, tabular inductive biases from tree-based models are inherited by the entire synthetic data generation model through the construction of the intermediate input sequence of tokens, thereby improving the quality of the generated synthetic data.
-Leaf indices are also provided as prompts during generation, allowing indirect yet informative hints to generate more realistic data.
+In tree $\mathcal{T}$, there is a leaf index matrix $\mathbf{J}=[j_{ik}]\in\mathbb{N}^{n\times T}$ of $\mathbf{X}$ where $j_{ik}\in\{1,2,\dots,l_{k}\}$ represents the leaf index of $i$-th row $\mathbf{X}_{i}$ in the $k$-th tree.
+Leaf indices provide flattened positional information of each row in a tree, and hence inherently introduce the non-smooth and non-rotationally invariant inductive biases of trees to capture the non-smoothness and potentially low-correlations of tabular data.
+With leaf indices prepended to actual data values, although the subsequent model is still a neural network, tabular-specific inductive biases from tree-based models are inherited by the entire generative model through the construction of the intermediate input token sequence (i.e., indicated by the red arrows in Fig.
+2), thereby improving the quality of the synthetic data.
+Leaf indices are also served as prompts during inference, allowing indirect yet informative hints to generate more realistic data.
 3.2 Tokenizer
-To model the multimodal distribution of continuous values and reduce tokenized value sizes to eventually reduce the model size, we devise a tailored data tokenizer for tabular data (“Tokenizer” in Fig.
+To model the multimodal distribution of continuous values and reduce the vocabulary size along with the tokenized input sequence length to eventually reduce the model size, we devise a tailored data tokenizer for tabular data (“Tokenizer” in Fig.
 2).
 For each continuous feature, we quantize the values so that it is of a more natural data format for transformers to learn.
-However, unlike TabMT (Gulati & Roysdon, 2024) that uses a single quantizer, we use a dual-quantizer (i.e., double quantizers) to capture multimodal continuous distribution and describe the precise values respectively.
-Transformers do not expect near-Gaussian-distributed input tokens, so we capture the multimodal distribution simply by a K-Means quantizer (MacQueen, 1967) with a small number of clusters (e.g.
-$K=10$).
-Following that, the original value is then quantized into a large number of quantiles (e.g.
-$Q=1000$), based on which original values are recovered, and thus the values are highly precise.
-By the dual-quantization tokenization method, each numeric value are encoded into 2 discrete values: a bin ID in the K-Means quantizer, and a quantile ID in the quantile quantizer, making 2 tokens sufficient to represent a numeric value with useful and sufficient information.
+However, unlike TabMT (Gulati & Roysdon, 2023) that uses a single quantizer, we use a dual-quantizer (i.e., double quantizers) to capture multimodal continuous distribution and describe the precise values respectively.
+Transformers do not expect near-Gaussian-distributed input tokens, so we capture the multimodal distribution simply by a K-Means quantizer (MacQueen, 1967) with a small number of clusters (e.g., $K=10$).
+Following that, the original value is then quantized into a large number of quantiles (e.g., $Q=1000$), based on which original values are recovered, and thus the values are highly precise.
+Employing the dual-quantization tokenization scheme, each numeric value is encoded into two discrete values: a bin ID in the K-Means quantizer and a quantile ID in the quantile quantizer, allowing two tokens to represent a numeric value with valuable and sufficient information.
 Categorical features are simply tokenized by label encoding, requiring 1 token.
@@
 Let $c_{i}\in\mathbb{N},b_{i}\in\{0,1,\dots,K\},q_{i}\in\{0,1,\dots,Q\}$ denote the number of categories, K-Means bins, and quantiles per column respectively, where $c_{i}=0$ in continuous features and $b_{i}=q_{i}=0$ in categorical features.
-Then, the domain for a categorical value is $\{1,2,\dots,c_{i}\}\subseteq\mathbb{N}_{+}$, and the domain for a continuous value is $\{1,2,\dots,b_{i}\}\times\{1,2,\dots,q_{i}\}\subseteq\mathbb{N}_{+}^{2}$.
+Then, the domain for a categorical value is $\{1,2,\dots,c_{i}\}\subseteq\mathbb{N}^{+}$, and the domain for a continuous value is $\{1,2,\dots,b_{i}\}\times\{1,2,\dots,q_{i}\}\subseteq(\mathbb{N}^{+})^{2}$.
 3.3 Auto-regressive Transformer
@@
 The input token sequence is constructed by concatenating the outcome from the tree model and tokenizer.
-Totally 5 major types of tokens are required and the vocabulary size is $V=n_{l}+n_{c}+n_{b}+n_{q}+3$.
-The number of tokens in a constructed sequence of tokens corresponding to one row is $L=2+T+m_{d}+2m_{c}$, including BOS and EOS tokens.
-Both of them are much lower than the need for natural language models.
-A shared set of leaf tokens across all trees and similarly shared sets of categorical, bin, and quantile tokens across all features are used.
+It requires 5 types of tokens and a vocabulary size of $V=n_{l}+n_{c}+n_{b}+n_{q}+3$.
+The number of tokens in a sequence corresponding to one row is $L=2+T+m_{d}+2m_{c}$, including BOS and EOS tokens.
+Both the vocabulary size and sequence length are much smaller than the need for natural language models.
+A shared set of leaf tokens across all trees and similarly shared sets of category, bin, and quantile tokens across all features are used.
 Distinctions between different trees and features are effectively encoded through positional information, leveraging the positional embeddings inherent in most transformers.
@@
 Consequently, tokens can be sampled exclusively from this precomputed set of valid options during generation.
-This ensures that every generated token sequence corresponds to a valid row, eliminating the need for rejecting invalid samples – a common practice in existing methods (Borisov et al., 2023; Solatorio & Dupriez, 2023; Zhao et al., 2023) – and significantly accelerating the sampling process.
+This ensures that every generated token sequence corresponds to a valid row, eliminating the need for rejecting invalid sequences (Borisov et al., 2023; Solatorio & Dupriez, 2023; Zhao et al., 2023), thus significantly accelerating the sampling process.
 Additionally, transformers are prone to memorization, particularly when trained on tabular data, where datasets are much smaller compared to typical natural language corpora (Borisov et al., 2023).
-To avoid over-fitting in terms of memorization, we heavily mask the input and maintain the target output unmasked.
+To avoid memorization, we heavily mask the input and maintain the target output unmasked.
 The rest of the training and generation settings are similar to the training and generation of classical causal language models, except for a modified loss function to account for the ordinal relations between quantile tokens, with details introduced in Section 3.4.
@@
 To address this, we replace the vanilla CEL with a specialized ordinal CEL for quantile tokens.
-Formally, let the predicted logits for a token be $\mathbf{z}\in\mathbb{R}^{V}$, and let $t\in\{1,\dots,V\}$ be the target label. Recall the classical definition of CEL as given in Equation 1.
-$\mathcal{L}_{\text{ce}}(\mathbf{z},t)=-\log\frac{e^{z_{t}}}{\sum_{i=1}^{V}e^{z_{i}}}$ (1)
+Formally, let the predicted logits for a token be $\mathbf{z}\in\mathbb{R}^{V}$, the probability after softmax be $\mathbf{p}\in[0,1]^{V}$ ($\|\mathbf{p}\|_{1}=1$), and let $t\in\{1,\dots,V\}$ be the target label.
+Recall the classical definition of CEL as given in Equation 1.
+$\mathcal{L}_{\text{ce}}(\mathbf{z},t)=-\log\frac{e^{z_{t}}}{\sum_{i=1}^{V}e^{z_{i}}}=-\log p_{t}\vspace{-0.2em}$ (1)
 We define ordinal cross-entropy loss (OCEL) as a variant of CEL where each term is weighted.
@@
 Theorem 1.
-If $w_{ti}\geq 0$ is a function of the distance of the current class $i$ to the target class $t$, namely, $w_{ti}=f(|t-i|)$ where $f:\mathbb{N}\rightarrow\mathbb{R}_{+},|t-i|\mapsto w_{i}$, and $f$ is monotonically increasing, then the OCEL penalizes adjacent classes less than farther classes.
+Let $f(|t-i|)$ be a variant of the distance between current class $i$ and target class $t$ when $f$ is non-negative and monotonically increasing.
+Let the weighted sum of this variant of distances to target incurred by some logit $\mathbf{z}$ be $D(\mathbf{z})=\sum_{i=1}^{V}p_{i}w_{i}$.
+If $w_{ti}=f(|t-i|)$ for some $f$, then for $\mathbf{z},\mathbf{z}^{*}$ with $\mathcal{L}_{\text{ce}}(\mathbf{z},t)=\mathcal{L}_{\text{ce}}(\mathbf{z}^{*},t)$ and $D(\mathbf{z})<D(\mathbf{z}^{*})$, we must have $\mathcal{L}_{\text{oce}}(\mathbf{z},t)<\mathcal{L}_{\text{oce}}(\mathbf{z}^{*},t)$.
 Both Lemma 1 and Theorem 1 are intuitive, and we provide rigorous proofs in Appendix A.1-A.2.
-Theorem 1 requires the definition of $w_{ti}=f(|t-i|)$ that satisfies the monotonically increasing constraint to be used for OCEL.
-Any satisfying definition works technically.
+Theorem 1 implies that the OCEL penalizes (value being smaller) classes closer to the target class less than farther ones.
+Any definition of $w_{ti}=f(|t-i|)$ satisfying the monotonicity constraint works for OCEL technically.
 In this paper, we define the weight $w_{ti}$ for $z_{i}$ to be obtained by Equation 3.
@@
 where $\sigma=0.005$ is a scaling factor of the distance, and $m=0.5$ is the minimum weight (used on $i=t$). Detailed explanation and analysis of it can be found in Appendix A.3.
+Figure 3: Performance comparison between different models in terms of fidelity.
+The higher the values are, the better the fidelity is.
+Top-4 performing baselines by fidelity mean scores are visualized.
 Note that not all tokens are part of the ordinal relation, so the OCEL is applied only to valid quantile tokens at the corresponding positions.
 To enable the model to learn to generate valid quantile tokens, we introduce an additional modified CEL to distinguish valid tokens from invalid ones, which complements OCEL.
-Formally, let valid quantile tokens have IDs in the range $[Q_{s},Q_{e})$ The valid token group loss (for quantile tokens at this position) is defined by Equation 4 when $t\in[Q_{s},Q_{e})$.
+Formally, let valid quantile tokens have IDs in the range $[Q_{s},Q_{e})$.
+The valid token group loss (for quantile tokens at this position) is defined by Equation 4 when $t\in[Q_{s},Q_{e})$.
 $\mathcal{L}_{\text{vg}}(\mathbf{z},t)=-\log\frac{\sum_{i\in[Q_{s},Q_{e})}e^{z_{i}}}{\sum_{i=1}^{V}e^{z_{i}}}\vspace{-0.5em}$ (4)
@@
 4 Experiments
+Figure 4: Comparison of marginal densities for generated data from baseline tabular auto-regressive transformers and TabTreeFormer against real data on representative multimodal continuous columns.
+All models use Distill-GPT2 as the backbone.
 4.1 Experimental Setup
@@
 Baseline models.
-We compare the performance of TabTreeFormer against several influential and high-performing methods in tabular data generation including non-neural networks, GANs, VAEs, diffusion models, and auto-regressive transformers: ARF (Watson et al., 2023), CTAB-GAN+ (CTAB+) (Zhao et al., 2024), CTGAN (Xu et al., 2019), TVAE (Xu et al., 2019), TabDDPM (Kotelnikov et al., 2023), TabSyn (Zhang et al., 2024), and GReaT (Borisov et al., 2023), REaLTabFormer (RTF) (Solatorio & Dupriez, 2023).
+We compare the performance of TabTreeFormer against several influential and high-performing methods in tabular data generation including non-neural networks, GANs, VAEs, diffusion models, and auto-regressive transformers: ARF (Watson et al., 2023), CTAB-GAN+ (CTAB+) (Zhao et al., 2024), CTGAN (Xu et al., 2019), TVAE (Xu et al., 2019), TabDDPM (Kotelnikov et al., 2023), TabSyn (Zhang et al., 2024), GReaT (Borisov et al., 2023) and REaLTabFormer (RTF) (Solatorio & Dupriez, 2023).
 Implementation of baselines is described in Appendix C.3.
+Figure 5: Performance comparison of TabTreeFormer versus top-4 performing baselines in fidelity on pair-wise correlation in 3 experimented datasets.
+Real (absolute) correlation values are presented on the left, and the absolute error in correlation values in synthetic data from different models, capped at 0.4 for better visibility, are shown at the right (the darker, the better).
+The left-top features are categorical features, and the right-bottom features are numeric features.
+Figure 6: Real vs. synthetic abs. corr.
+Figure 7: Feature count vs. Trend score improvement of TabTreeFormer from baselines.
+Table 2: MLE performance between different models.
+The top two rows calculate the global average scores and relative error (w.r.t real data).
+The middle three rows show the average performance in each ML model, and the bottom two rows show the average performances in classification (clf.) and regression (reg.).
+Classification tasks are evaluated by weighted AUC ROC scores, and regression tasks are evaluated by $R^{2}$ scores.
+The best scores and the second best scores are highlighted in bold with and without underscore, respectively.
+ARF CTAB+ CTGAN TabDDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
+Avg.
+($\uparrow$) $0.843_{\pm 0.155}$ $0.680_{\pm 0.271}$ $0.810_{\pm 0.189}$ $0.552_{\pm 0.865}$ $0.800_{\pm 0.206}$ $0.866_{\pm 0.130}$ $0.865_{\pm 0.123}$ $0.813_{\pm 0.189}$ $0.863_{\pm 0.127}$ $\boldsymbol{0.878}_{\pm 0.113}$ $\underline{\boldsymbol{0.880}}_{\pm 0.111}$
+RE ($\downarrow$) $0.057_{\pm 0.077}$ $0.244_{\pm 0.274}$ $0.099_{\pm 0.137}$ $0.483_{\pm 1.214}$ $0.110_{\pm 0.170}$ $0.027_{\pm 0.040}$ $0.027_{\pm 0.029}$ $0.095_{\pm 0.142}$ $0.029_{\pm 0.040}$ $\boldsymbol{0.017}_{\pm 0.023}$ $\underline{\boldsymbol{0.016}}_{\pm 0.023}$
+LN ($\uparrow$) $0.835_{\pm 0.176}$ $0.686_{\pm 0.258}$ $0.810_{\pm 0.193}$ $0.570_{\pm 0.882}$ $0.800_{\pm 0.197}$ $0.851_{\pm 0.153}$ $0.849_{\pm 0.156}$ $0.814_{\pm 0.163}$ $0.853_{\pm 0.154}$ $\boldsymbol{0.861}_{\pm 0.145}$ $\underline{\boldsymbol{0.862}}_{\pm 0.143}$
+RF ($\uparrow$) $0.842_{\pm 0.156}$ $0.681_{\pm 0.273}$ $0.806_{\pm 0.200}$ $0.566_{\pm 0.910}$ $0.798_{\pm 0.220}$ $0.873_{\pm 0.122}$ $0.872_{\pm 0.109}$ $0.810_{\pm 0.210}$ $0.866_{\pm 0.121}$ $\boldsymbol{0.888}_{\pm 0.095}$ $\underline{\boldsymbol{0.891}}_{\pm 0.090}$
+XGB ($\uparrow$) $0.851_{\pm 0.147}$ $0.672_{\pm 0.307}$ $0.814_{\pm 0.194}$ $0.519_{\pm 0.897}$ $0.802_{\pm 0.222}$ $0.874_{\pm 0.125}$ $0.874_{\pm 0.110}$ $0.814_{\pm 0.209}$ $0.870_{\pm 0.116}$ $\boldsymbol{0.884}_{\pm 0.102}$ $\underline{\boldsymbol{0.886}}_{\pm 0.101}$
+clf.($\uparrow$) $0.907_{\pm 0.074}$ $0.750_{\pm 0.227}$ $0.895_{\pm 0.065}$ $0.850_{\pm 0.114}$ $0.868_{\pm 0.130}$ $0.917_{\pm 0.064}$ $0.913_{\pm 0.071}$ $0.890_{\pm 0.081}$ $0.918_{\pm 0.062}$ $\underline{\boldsymbol{0.922}}_{\pm 0.063}$ $\boldsymbol{0.922}_{\pm 0.065}$
+reg.
+($\uparrow$) $0.586_{\pm 0.122}$ $0.400_{\pm 0.266}$ $0.472_{\pm 0.127}$ $-0.641_{\pm 1.467}$ $0.528_{\pm 0.239}$ $0.660_{\pm 0.122}$ $0.674_{\pm 0.095}$ $0.506_{\pm 0.187}$ $0.643_{\pm 0.061}$ $\boldsymbol{0.698}_{\pm 0.083}$ $\underline{\boldsymbol{0.710}}_{\pm 0.092}$
 Metrics.
@@
 We evaluate based on the established Train-on-Synthetic, Test-on-Real (TSTR) machine learning efficacy (MLE) evaluation framework (Xu et al., 2019), namely, synthetic data generated based on real training data is tested on a hold-out real test set.
-Three models are used: linear (i.e.
-linear and logistic regression; LN), random forest (RF) (Breiman, 2001), and XGBoost (XGB) (Chen & Guestrin, 2016).
-Better generative models tend to yield better scores on downstream tasks.
+Three models are used: linear (i.e., linear and logistic regression; LN), random forest (RF) (Breiman, 2001), and XGBoost (XGB) (Chen & Guestrin, 2016).
+Classification performance is evaluated by weighted AUC ROC and regression is evaluated by $R^{2}$.
+For both metrics, better generative models tend to yield higher scores on downstream tasks.
   • Privacy is a crucial criterion when the synthetic data serves the use case of privacy-preserving data sharing.
@@
 We compare the DCR from synthetically generated data and from hold-out real (test) data to the real training data and privacy preservation is demonstrated by DCRs calculated on the latter being no smaller than the former, tested by Mann-Whitney U Test (Mann & Whitney, 1947).
-Privacy is at risk of disclosure when $p$-values are smaller than 0.05.
+Privacy is at risk when $p$-values are smaller than 0.05.
   • Efficiency showcases the model sizes and computation time for training and generation of different models. Under comparable performance, smaller models and faster computation are favored.
 4.2 Fidelity
-Figure 3: Performance comparison between different models in terms of fidelity.
-The higher the values are, the better the fidelity is.
-Top-4 performing baselines by fidelity mean scores are visualized.
-Figure 4: Visualization of the marginal densities of the generated data of baseline tabular auto-regressive transformers and TabTreeFormer in comparison to the real data on some representative multimodal-distributed continuous columns.
-All these models use Distill-GPT2 (or its structure and hyperparameters) as backbone.
-Figure 5: Performance comparison of TabTreeFormer versus top-4 performing baselines in fidelity on pair-wise correlation in 3 experimented datasets.
-Real (absolute) correlation values are presented on the left, and the absolute error in correlation values in synthetic data from different models, capped at 0.4 for better visibility, are shown at right (the darker, the better).
-The left-top features are categorical fatures, and right-bottom features are numeric features.
-Figure 6: Real vs. synthetic abs. corr.
-Figure 7: Feature count vs. “Trend” score improvement of TabTreeFormer from some baselines.
 Marginal density distribution of columns (Shape).
-TabTreeFormer achieves comparable performance in “Shape” metric, as shown in Fig.
+TabTreeFormer achieves comparable performance to baseline models in “Shape” metric, as shown in Fig.
 3.
@@
 7.
-This also demonstrates the advantage of introducing the inductive bias on low-correlations so that the models learn relations among a larger number of features better by filtering out less correlated features.
+This also demonstrates the advantage of introducing the inductive bias to address low-correlations so that the models learn relations among a larger number of features better by filtering out less correlated features.
 4.3 Utility (MLE)
-Table 2: MLE performance between different models.
-The top two rows calculate the global average scores and relative error (w.r.t real data).
-The middle three rows show the average performance in each ML model, and the bottom two rows show the average performances in classification (clf.) and regression (reg.).
-Classification tasks are evaluated by weighted AUC ROC scores, and regression tasks are evaluated by $R^{2}$ scores.
-Best scores and second best scores are highlighted in bold with and without underscore, respectively.
-ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
-Avg.
-$0.843_{\pm 0.155}$ $0.680_{\pm 0.271}$ $0.810_{\pm 0.189}$ $0.552_{\pm 0.865}$ $0.800_{\pm 0.206}$ $0.866_{\pm 0.130}$ $0.865_{\pm 0.123}$ $0.813_{\pm 0.189}$ $0.863_{\pm 0.127}$ $\boldsymbol{0.878}_{\pm 0.113}$ $\underline{\boldsymbol{0.880}}_{\pm 0.111}$
-RE $0.057_{\pm 0.077}$ $0.244_{\pm 0.274}$ $0.099_{\pm 0.137}$ $0.483_{\pm 1.214}$ $0.110_{\pm 0.170}$ $0.027_{\pm 0.040}$ $0.027_{\pm 0.029}$ $0.095_{\pm 0.142}$ $0.029_{\pm 0.040}$ $\boldsymbol{0.017}_{\pm 0.023}$ $\underline{\boldsymbol{0.016}}_{\pm 0.023}$
-LN $0.835_{\pm 0.176}$ $0.686_{\pm 0.258}$ $0.810_{\pm 0.193}$ $0.570_{\pm 0.882}$ $0.800_{\pm 0.197}$ $0.851_{\pm 0.153}$ $0.849_{\pm 0.156}$ $0.814_{\pm 0.163}$ $0.853_{\pm 0.154}$ $\boldsymbol{0.861}_{\pm 0.145}$ $\underline{\boldsymbol{0.862}}_{\pm 0.143}$
-RF $0.842_{\pm 0.156}$ $0.681_{\pm 0.273}$ $0.806_{\pm 0.200}$ $0.566_{\pm 0.910}$ $0.798_{\pm 0.220}$ $0.873_{\pm 0.122}$ $0.872_{\pm 0.109}$ $0.810_{\pm 0.210}$ $0.866_{\pm 0.121}$ $\boldsymbol{0.888}_{\pm 0.095}$ $\underline{\boldsymbol{0.891}}_{\pm 0.090}$
-XGB $0.851_{\pm 0.147}$ $0.672_{\pm 0.307}$ $0.814_{\pm 0.194}$ $0.519_{\pm 0.897}$ $0.802_{\pm 0.222}$ $0.874_{\pm 0.125}$ $0.874_{\pm 0.110}$ $0.814_{\pm 0.209}$ $0.870_{\pm 0.116}$ $\boldsymbol{0.884}_{\pm 0.102}$ $\underline{\boldsymbol{0.886}}_{\pm 0.101}$
-clf.
-$0.907_{\pm 0.074}$ $0.750_{\pm 0.227}$ $0.895_{\pm 0.065}$ $0.850_{\pm 0.114}$ $0.868_{\pm 0.130}$ $0.917_{\pm 0.064}$ $0.913_{\pm 0.071}$ $0.890_{\pm 0.081}$ $0.918_{\pm 0.062}$ $\underline{\boldsymbol{0.922}}_{\pm 0.063}$ $\boldsymbol{0.922}_{\pm 0.065}$
-reg.
-$0.586_{\pm 0.122}$ $0.400_{\pm 0.266}$ $0.472_{\pm 0.127}$ $-0.641_{\pm 1.467}$ $0.528_{\pm 0.239}$ $0.660_{\pm 0.122}$ $0.674_{\pm 0.095}$ $0.506_{\pm 0.187}$ $0.643_{\pm 0.061}$ $\boldsymbol{0.698}_{\pm 0.083}$ $\underline{\boldsymbol{0.710}}_{\pm 0.092}$
-The MLE results are presented in Table 2, whose raw scores are in Table 10 in Appendix D.2, show that TabTreeFormer outperforms baseline models steadily regardless of the complexity of the ML model used, task, and dataset.
-TabTreeFormer-S achieves similar performance as the best-performing baselines (REaLTabForme and TabSyn), and the M and L versions of TabTreeFormer outperforms them by a significant margin of around 40%, assessed by relative errors of ML performance compared to real data.
-Furthermore, out of the 30 MLE scores, TabTreeFormer-M or -L outperforms all other baselines in 20 cases (2/3), and they both outperform all other baselines in 16 cases ($\geq$1/2).
-As an important foundation of the superior capability of tree-based models in capturing potentially low-correlations inductive bias, rotational invariance of models is an aspect worthwhile analyzing.
+The MLE results presented in Table 2 show that TabTreeFormer outperforms baseline models steadily regardless of the complexity of task, dataset, and ML model used.
+TTF-S achieves similar performance as the best-performing baselines (RTF and TabSyn), and TTF-M and TTF-L outperform them by a large margin of almost 40% in terms of relative errors in ML performance compared to real data.
+More MLE results can be found in Table 10 in Appendix D.2.
+As an important foundation of tree-based models’ superior capability in capturing potentially low-correlations, rotational invariance of models is an aspect particularly worthwhile analyzing.
 As visualized in Fig.
-9, rotationally invariant generators (e.g.
-CTAB-GAN+, CTGAN, GReaT, TVAE) are generally worse than non-rotationally invariant counterparts (e.g.
-REaLTabFormer and our TabTreeFormer by positional embedding in a fixed order and TabSyn by VAE-encoding of data).
-Also, TabTreeFormer’s design outperforms rotationally invariant baseline generators more on better-performing non-rotationally invariant downstream models (RF and XGB) than on rotationally invariant downstream models (LN).
-This again emphasizes the importance of applying non-rotationally invariant models to introduce tabular inductive bias (for low correlations).
+9, rotationally invariant generators (e.g., CTAB-GAN+, CTGAN, GReaT, TVAE) are generally worse than non-rotationally invariant counterparts (e.g., REaLTabFormer and TabTreeFormer with a positional embedding in fixed order and TabSyn with VAE-encoding of data).
+Also, the advantage of TabTreeFormer over rotationally invariant baseline generators is more manifest on better-performing non-rotationally invariant downstream models (RF and XGB) than on rotationally invariant ones (LN).
+This again emphasizes the importance of applying non-rotationally invariant models to introduce inductive bias for low correlations.
 Figure 8: MLE relative error to real data with TabTreeFormer-L. On the left of the red line are rotationally invariant models.
@@
 4.4 Privacy (DCR)
-Table 3: Summarized performance comparison between different models in terms of privacy.
-The first row shows the number of violations, and the second row shows the minimum $p$ value for privacy leakage risk, the closer the value to 0, the higher the risk of privacy leakage.
+Table 3: Privacy comparison of models.
+The first row shows the number of violations, and the second row shows the minimum $p$ value for privacy leakage risk.
+The closer the value to 0, the higher the risk of privacy leakage.
 Risky values are highlighted in red.
@@
 min. $p$ 0.071 0.000 0.097 0.141 0.003 0.045 0.061 0.111
-As summarized in Table 3 (raw scores in Table 11 in Appendix D.3), TabTreeFormer-M and L are safe in all the 10 datasets.
-TabTreeFormer-S has a privacy leakage in only one dataset, with the $p$-value very close to 0.05.
-In comparison, GReaT has severe privacy concerns in 5 out of 10 datasets.
-Therefore, while auto-regressive transformers have privacy leakage concerns, TabTreeFormer and REaLTabFormer are generally privacy-preserving, validating the effectiveness of using masking in preserving privacy for tabular auto-regressive transformers.
-4.5 Efficiency
 Table 4: Relative performance change of TabTreeFormer-M compared to TabTreeFormer-L.
@@
 Trend $-0.436\%$ Avg. MLE relative error $-9.624\%$
-Observing results reported above, the performance of TabTreeFormer-M is better than all baselines in most metrics, while it has only about 5M parameters and its most competitive counterparts have more than 20M (TabSyn) or even more than 80M parameters (GReaT and REaLTabFormer).
-Also, as shown in Table 4, there is no significant performance difference between TabTreeFormer-M and TabTreeFormer-L, although their sizes differ by a factor of 8.
+As summarized in Table 3TTF-M and TTF-L are privacy-resilient in all the 10 datasets.
+TabTreeFormer-S has a privacy leakage in only one dataset, with the $p$-value very close to 0.05.
+In comparison, GReaT demonstrates severe privacy concerns in 5 out of 10 datasets.
+Although auto-regressive transformers suffer from privacy leakage concerns, TabTreeFormer and REaLTabFormer are generally privacy-preserving, validating the effectiveness of token masking in protecting privacy for tabular auto-regressive transformers.
+Raw $p$-values per experiment are presented in Table 11 in Appendix D.3.
+4.5 Efficiency
+The efficiency of TabTreeFormer is mainly demonstrated by its ability to achieve better performance with a smaller model size.
+TTF-M surpasses all baselines in most metrics (i.e., fidelity, MLE, and DCR) as shown in Section 4.2-4.4.
+Notably, it only has approximately 5M parameters while the most competitive counterparts have more than 20M (TabSyn) or even more than 80M parameters (GReaT and REaLTabFormer).
+Also, as shown in Table 4, there is no significant performance gap between TTF-M and TTF-L, although their sizes differ eight times.
 This demonstrates the capability of TabTreeFormer in maintaining high efficiency without significant compromise in effectiveness.
-Table 5: Training time (s) comparison of TabTreeFormer compared to other auto-regressive transformer baselines on large datasets.
-Best values are highlighted in bold with underscore, and second best values are highlighted in bold without underscore.
+Table 5: Training time (s) comparison of TabTreeFormer and other auto-regressive transformer baselines on large datasets.
+The best values are highlighted in bold with an underscore, and the second best values are highlighted in bold without underscore.
 Dataset adult bank cahouse
@@
 TTF-L $\boldsymbol{606.790}_{\pm 277.828}$ $\boldsymbol{744.726}_{\pm 221.487}$ $946.903_{\pm 102.584}$
-Compared to baseline auto-regressive transformers, TabTreeFormer demonstrates significant advantage in terms of efficiency, especially in training time on large datasets (see Table 5) and in generation time on datasets with a large number of features (see Fig.
+Meanwhile, compared to baseline auto-regressive transformers, TabTreeFormer demonstrates a significant advantage in terms of training time on large datasets (see Table 5) and generation time on datasets with a large number of features (see Fig.
 9), which is consistent with our expectation due to the smaller model size and sampling without rejection.
@@
 4.6 Ablation Study
-Table 6: Ablation studies. The values show the relative score change when certain component in TabTreeFormer-S is taken out.
+Table 6: Design component ablation. Relative score change when the corresponding component in TTF-S is taken out.
 LN RF XGB
@@
 w/o OCEL $+0.452\%$ $-0.967\%$ $-1.446\%$
-On each core components of TabTreeFormer, including tree-based model (TBM), dual-quantization tokenization (DQT), and OCEL, we apply ablation study using TabTreeFormer-S on 2 randomly selected datasets: credit and diabetes, evaluated using MLE.
-The results, presented in Table 6 with raw values presented in Table 14 in Appendix D.6, verify the usefulness of all the core components.
+We study the necessity of each building block in TabTreeFormer.
+It can be seen from Table 6 that TabTreeFormer experiences different degrees of noticeable performance degradation when we eliminate the tree-based model (TBM), dual-quantization tokenization (DQT), and OCEL, respectively, which verifies the indispensability of all building blocks.
+More results are presented in Appendix D.6.
 5 Conclusion
-In this paper, we propose TabTreeFormer, a hybrid tree-transformer for generating high-quality tabular data.
-TabTreeFormer inherits unique inductive biases from a tree-based model and also captures multimodal distribution to optimize the data representations of numeric values via a dual-quantization tokenizer.
-We comprehensively evaluate the performance of TabTreeFormer and show that it demonstrates superior fidelity, utility, privacy and outstanding efficiency with significantly smaller model size and computation time.
+In this paper, we propose TabTreeFormer, a hybrid tree-transformer for high-quality tabular data generation.
+TabTreeFormer inherits tabular-specific inductive biases from a tree-based model and captures multimodal distribution for optimized numeric data representations via a dual-quantization tokenizer.
+We comprehensively evaluate the performance of TabTreeFormer and show that it demonstrates outstanding fidelity, utility, privacy, and efficiency with significantly smaller model sizes and computation time.
+Acknolwedgements
+We would like to acknowledge Rishav Chourasia for helpful discussions.
 Appendix A Supplementary Loss Function Explanation and Analysis
@@
 Proof.
-Suppose we have $\mathbf{z}$ and $\mathbf{z}^{*}$ as logits on vocabulary size $V$ and target label $t$, which satisfy both the following constraints:
-  • For some $i\neq j\in\{1,\dots,V\}\setminus$ and $|t-i|\neq|t-j|$, $a=z_{i}=z_{j}^{*}\neq z_{j}=z_{i}^{*}=b$ for some $a,b\in\mathbb{R}$.
-  • $z_{k}=z_{k}^{*},\forall k\in\{0,1,\dots,V-1\}\setminus\{i,j,t\}$.
-Without loss of generality, suppose $d_{1}=|t-i|<|t-j|=d_{2},a<b$. Note that $w_{k}$ is monotonically increasing with $|t-k|$.
-Thus, $e^{a}<e^{b},w_{ti}<w_{tj}$.
-This gives $\left(w_{ti}-w_{tj}\right)\left(e^{a}-e^{b}\right)<0$, so $P=w_{ti}e^{z_{i}}+w_{tj}e^{z_{j}}=w_{ti}e^{a}+w_{tj}e^{b}<w_{ti}e^{z_{i}^{*}}+w_{tj}e^{z_{j}^{*}}=w_{ti}e^{b}+w_{tj}e^{a}=Q$ for some $P,Q\in\mathbb{R}$.
-Substitute into $\mathcal{L}_{\text{oce}}(\mathbf{z},t)$ and $\mathcal{L}_{\text{oce}}(\mathbf{z}^{*},t)$,
-$\mathcal{L}_{\text{oce}}(\mathbf{z},t)$ $=-\log\frac{w_{tt}e^{z_{t}}}{P+\sum_{k\neq i,j}w_{k}e^{z_{k}}}$
-$\mathcal{L}_{\text{oce}}(\mathbf{z}^{*},t)$ $=-\log\frac{w_{tt}e^{z_{t}}}{Q+\sum_{k\neq i,j}w_{k}e^{z_{k}}}$
-Hence, $\mathcal{L}_{\text{oce}}(\mathbf{z},t)<\mathcal{L}_{\text{oce}}(\mathbf{z}^{*},t)$. The logits with higher logit value at position closer to target token is penalized less. ∎
+Express $\mathcal{L}_{\text{oce}}$ and $\mathcal{L}_{\text{ce}}$ using $\mathbf{p}$, we have
+$\mathcal{L}_{\text{ce}}(\mathbf{z},t)$ $=-\log p_{t}$ (6)
+$\mathcal{L}_{\text{oce}}(\mathbf{z},t)$ $=-\log\frac{w_{tt}e^{z_{t}}}{\sum_{i=1}^{V}w_{ti}e^{z_{i}}}$
+$=-\log\frac{w_{tt}\cdot\frac{e^{z_{t}}}{\sum_{i=1}^{V}e^{z_{i}}}}{\sum_{i=1}^{V}w_{ti}\cdot\frac{e^{z_{i}}}{\sum_{j=1}^{V}e^{z_{j}}}}$
+$=-\log\frac{w_{tt}p_{t}}{\sum_{i=1}^{V}w_{ti}p_{i}}$
+$=-\log\frac{w_{tt}p_{t}}{D(\mathbf{z})}$ (7)
+By $\mathcal{L}_{\text{ce}}(\mathbf{z},t)=\mathcal{L}_{\text{ce}}(\mathbf{z}^{*},t)$, we have $p_{t}=p_{t}^{*}$.
+Given $D(\mathbf{z})<D(\mathbf{z}^{*})$, we must have $\mathcal{L}_{\text{oce}}(\mathbf{z},t)<\mathcal{L}_{\text{oce}}(\mathbf{z}^{*},t)$.
 A.3 Detailed Explanations and Analysis of the Weight Function (Equation 3)
@@
 TabTreeFormer S M L
-Approx. #Params 1 5 40
+Approx. #Params (in M) 1 5 40
 hidden state dimensions 128 256 768
@@
 C.3 Baseline Implementation
-We use the Synthcity (Qian et al., 2024) implementation for ARF (Watson et al., 2023), CTGAN (Xu et al., 2019), GReaT (Borisov et al., 2023), TabDDPM (Kotelnikov et al., 2023), and TVAE (Xu et al., 2019).
+We use the Synthcity (Qian et al., 2023) implementation for ARF (Watson et al., 2023), CTGAN (Xu et al., 2019), GReaT (Borisov et al., 2023), TabDDPM (Kotelnikov et al., 2023), and TVAE (Xu et al., 2019).
 We also include CTAB-GAN+ (Zhao et al., 2024), REaLTabFormer (Solatorio & Dupriez, 2023), and TabSyn (Zhang et al., 2024) using their official code on GitHub or public SDK.
@@
 The closer the values are to 1, the better the fidelity is.
-Best scores and second best scores are highlighted in bold with and without underscore, respectively.
+The best scores and the second best scores are highlighted in bold with and without underscore, respectively.
 Equal values in the first 3 digits may be compared by the 4-th or 5-th digits.
-Dataset Metric ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
+Dataset Metric ARF CTAB+ CTGAN TabDDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
 adult Shape $0.942_{\pm 0.001}$ $0.964_{\pm 0.005}$ $0.881_{\pm 0.008}$ $\underline{\boldsymbol{0.983}}_{\pm 0.001}$ $0.929_{\pm 0.000}$ $0.962_{\pm 0.001}$ $\boldsymbol{0.977}_{\pm 0.011}$ $0.877_{\pm 0.004}$ $0.946_{\pm 0.002}$ $0.954_{\pm 0.001}$ $0.956_{\pm 0.001}$
@@
 D.2 Supplementary Utility Results
-The raw experiment results for MLE is shown in Table 10. The results are also visualized in Figure 12.
+The raw experiment results for MLE is shown in Table 10.
+The results are also visualized in Figure 12.Out of the 30 MLE scores, TabTreeFormer-M or -L outperforms all other baselines in 20 cases (2/3), and they both outperform all other baselines in 16 cases ($\geq$1/2).
 Table 10: Performance comparison between different models in terms of MLE.
@@
 The last row shows the average relative error versus real.
-Best scores and second best scores are highlighted in bold with and without underscore, respectively.
+The best scores and the second best scores are highlighted in bold with and without underscore, respectively.
 Equal values in the first 3 digits may be compared by the 4-th or 5-th digits.
-Dataset ML real ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
+Dataset ML real ARF CTAB+ CTGAN TabDDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
 adult LN 0.914 $0.907_{\pm 0.000}$ $0.904_{\pm 0.001}$ $0.886_{\pm 0.009}$ $0.908_{\pm 0.000}$ $\boldsymbol{0.911}_{\pm 0.000}$ $\underline{\boldsymbol{0.911}}_{\pm 0.000}$ $0.910_{\pm 0.002}$ $0.893_{\pm 0.001}$ $0.905_{\pm 0.001}$ $0.908_{\pm 0.002}$ $0.910_{\pm 0.000}$
@@
 Table 11: Performance comparison between different models in terms of DCR. $p$-values less than $0.05$ is highlighted in red, indicating a high risk of privacy leakage.
-Dataset ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
+Dataset ARF CTAB+ CTGAN TabDDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
 adult $0.486_{\pm 0.077}$ $0.691_{\pm 0.424}$ $0.687_{\pm 0.272}$ $0.440_{\pm 0.125}$ $0.000_{\pm 0.000}$ $0.139_{\pm 0.131}$ $0.654_{\pm 0.149}$ $0.003_{\pm 0.004}$ $0.999_{\pm 0.001}$ $0.992_{\pm 0.011}$ $0.993_{\pm 0.006}$
@@
 AUC-ROC scores are reported, the smaller values mean that the synthetic records are harder to be detected.
-Best scores and second best scores are highlighted in bold with and without underscore, respectively.
+The best scores and the second best scores are highlighted in bold with and without underscore, respectively.
 Equal values in the first 3 digits may be compared by the 4-th or 5-th digits.
-Dataset ML ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
+Dataset ML ARF CTAB+ CTGAN TabDDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
 adult LN $0.794_{\pm 0.006}$ $0.632_{\pm 0.016}$ $0.946_{\pm 0.005}$ $\underline{\boldsymbol{0.585}}_{\pm 0.005}$ $0.714_{\pm 0.000}$ $\boldsymbol{0.602}_{\pm 0.000}$ $0.614_{\pm 0.031}$ $0.945_{\pm 0.003}$ $0.852_{\pm 0.001}$ $0.836_{\pm 0.001}$ $0.830_{\pm 0.001}$
@@
 Table 13: Comparison between different models in terms of time taken for training and generation. The values are in unit of seconds.
-Dataset Action ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
+Dataset Action ARF CTAB+ CTGAN TabDDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
 adult Train $258.758_{\pm 55.030}$ $787.379_{\pm 3.334}$ $753.259_{\pm 104.080}$ $285.153_{\pm 0.410}$ $10397.040_{\pm 9.203}$ $1875.904_{\pm 34.663}$ $1199.823_{\pm 146.744}$ $1111.927_{\pm 2.790}$ $166.807_{\pm 31.955}$ $235.554_{\pm 71.542}$ $606.790_{\pm 277.828}$
@@
 TabDiff - - $0.237_{\pm 0.398}$ $0.784_{\pm 1.765}$ $0.133_{\pm 0.230}$ $0.068_{\pm 0.134}$ $0.201_{\pm 0.279}$ $0.058_{\pm 0.112}$
-There are two recent tabular data generation works, TabMT (Gulati & Roysdon, 2024) and TabDiff (Shi et al., 2024), that have not published their code by the time of the composition of this paper.
+There are two recent tabular data generation works, TabMT (Gulati & Roysdon, 2023) and TabDiff (Shi et al., 2024), that have not published their code by the time of the composition of this paper.
 Unfortunately, although all papers report MLE scores, which can be a basis of comparison, the values from these papers cannot be directly taken for comparison because of the different settings of evaluation, including datasets, splits, ML model, and evaluation metric.